### v1

# Fast Real Evaluation Through Sound Mixed-Precision Tuning

Abstract.
Evaluating a real-valued expression to high precision is a key building block in computational mathematics, physics, and numerics. A typical implementation uses a uniform precision for each operation, and doubles that precision until the real result can be bounded to some sufficiently narrow interval. However, this is wasteful: usually only a few operations really need to be performed at high precision, and the bulk of the expression could use much lower precision. Uniform precision can also waste iterations discovering the necessary precision and then still overestimate by up to a factor of two.
We propose to instead use mixed-precision interval arithmetic to evaluate real-valued expressions. A key challenge is deriving the mixed-precision assignment both soundly and quickly. To do so, we introduce a sound variation of error Taylor series and condition numbers, specialized to interval arithmetic, that can be evaluated with minimal overhead thanks to an “exponent trick”. Our implementation, Reval, achieves a speed-up of 1.25×1.25$\times$1.25 × compared to the state-of-the-art Sollya tool, with the speed-up increasing to 2.99×2.$9$ on the most difficult input points. An examination of the precisions used with and without precision tuning shows that the speed-up results come from quickly assigning lower precisions for the majority of operations.
Interval arithmetic, precision tuning, correct rounding, rounding error, error Taylor series, condition numbers, arbitrary precision
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
Many tasks in computational science require computing real expressions to high precision, including computational number theory (Hakert et al., 2024; Bailey, 2001), physics (Bailey and Borwein, 2015), geometry (Seel, 2024), and commutative algebra (Grayson and Stillman, [n. d.]). Moreover, computing real expressions to high precision is also a computational bottleneck in the Remez or Chebyshev approximation algorithms (Driscoll et al., 2014; Remez, 1934; Lim et al., 2021); in numerical compilers like Herbie (Panchekha et al., 2015; Flatt and Panchekha, 2023); in implementations of math library functions like $\exp$ and $\operatorname{pow}$ (Bugzilla, 2006; archive of the discontinued LLVM Phabricator instance, 2020); and even in simple desktop calculators such as the one available on Android (Boehm, 2020).
Computing real expressions to high precision is challenging because computer arithmetic is subject to rounding error. Real number evaluation thus typically uses high-precision interval arithmetic (Chevillard et al., 2010; Panchekha et al., 2015; Johansson, 2016), in particular a strategy known as Ziv’s onion-peeling method (Ziv, 1991). This method first evaluates the real expression using interval arithmetic and checks if the resulting interval is narrow enough for whatever purpose the user may have in mind. If it’s not, Ziv’s strategy repeatedly re-evaluate the same expressions at higher and higher precision until a narrow-enough output interval is achieved. A typical implementation uses a single, uniform precision for evaluating the whole expression and doubles that precision with every re-evaluation.
However, evaluating the whole expression at a uniform precision is wasteful and slow: it forces us to evaluate the whole expression at the precision needs of its highest-precision operation. A typical example might be evaluating $\sin(\pi\cdot x)+2$ for a very large number $x\approx 10^{100}$. To compute an answer to, say, 20 decimal places, the multiplication by $\pi$ must be performed at extremely high precision, producing at least 120 correct output digits. However, the rest of the computation doesn’t need nearly this many bits: the output of $\sin$ is 0 and doesn’t need any digits, and the addition needs to produce only 20 correct digits. Using a uniform precision for this example would waste time and memory. And when a high precision is needed, the uniform precision-doubling approach must perform many wasteful iterations in order to discover the necessary high precision, which it might overestimate by a factor of two. A better approach would tune the precision of each operation individually.
We implement just that in Reval, a new library for real evaluation using intervals. [^1] The name is also a pun on Rival, the underlying interval library Reval uses. Reval uses a precision assignment algorithm to determine what precision to use for every operation in the program. The computed precision assignment is based on sound worst-case error bounds and can be computed quickly, with minimal overhead, and results in dramatically lower precision needs for the overwhelming majority of operations in an expression. This allows Reval to evaluate real expressions to high precision faster than state-of-the-art libraries such as Sollya (Chevillard et al., 2010).
Reval derives a sound floating-point error bound using an algorithm inspired by error Taylor series and condition numbers as used in FPTaylor (Solovyev et al., 2018), Satire (Das et al., 2020), and ATOMU (Zou et al., 2019). However, Reval’s error bounds are specialized to interval arithmetic, which allows for sound bounds without the need for a “higher-order error” term that requires second derivatives. Because Reval’s bounds only require first derivatives, they can be computed in time linear to the size of the expression, instead of the quadratic time required for “higher-order error” terms. To avoid slow arbitrary-precision computations, Reval uses an “exponent trick” which allows it to over-approximate the error bound using only operations over machine integers. The resulting error bounds can then be used for precision assignment, yielding sound, reasonably tight error bounds computed with minimal overhead.
We evaluate Reval on 4,359,168 computations (528 benchmarks each evaluated on 8 256 inputs) drawn from the Herbie 2.0 benchmark suite, and compare it to the state-of-the-art Sollya tool (Chevillard et al., 2010). Reval is 1.25×1.25$\times$1.25 × faster while successfully evaluating a similar number of computations, with an even larger speedup of 2.99×2.$9$ on the most difficult computations. A comparison with uniform precision-doubling shows that Reval’s performance advantage is due to executing most operations at substantially lower precision. Besides these quantitative measures, Reval’s precision assignment algorithm is also more robust to invalid inputs.
In summary, this paper contributes:
  • A sound error bound algorithm specialized to interval arithmetic (Section 4).
  • An “exponent trick” for computing these sound error bounds efficiently (Section 5).
  • Real expression evaluation using these error bounds for precision assignment (Section 6).
2. Overview
[…]
For this reason, real evaluation is typically done using interval arithmetic, which can certify that a computed answer is actually correct. With interval arithmetic, each value is represented by an interval $[x_{1},x_{2}]$ that contains the correct real-valued result. With an interval arithmetic library at, say, 83 bits, the interval for $\cos(x)$ is:
[0.99999999999999999999999948,0.99999999999999999999999959].0.999999999999999999999999480.99999999999999999999999959[0.99999999999999999999999948,0.99999999999999999999999959].[ 0.99999999999999999999999948 , 0.99999999999999999999999959 ] .
When subtracted from $1.0$, the resulting interval is approximately $[4.13\cdot 10^{-25},5.17\cdot 10^{-25}]$, and the final interval is $[4.13\cdot 10^{-13},5.17\cdot 10^{-13}]$. Note that this final interval includes the correct result—but is too wide to actually identify it. A higher precision, such as 160 bits, will compute a narrower interval like
[…]
and for this interval both endpoints round to the same double-precision number, $5\cdot 10^{-13}$. This proves that that is the correctly-rounded double-precision result.
Figure 1. In the standard uniform-precision approach (left), all operations share the same precision, which is doubled until a correct rounding is found. In Reval (right), precision tuning is performed if a correct rounding is not found, deriving a different precision for each operation.
This overall strategy—evaluating in arbitrary-precision interval arithmetic, and raising the precision until the interval is narrow enough—is the standard approach to accurately evaluating real expressions, diagrammed in Figure 1. The key challenge is choosing the precision. If the goal is 53 correct bits—a correct double-precision result—trying a uniform 80- and 160-bit evaluation is wasteful. For example, 160 bits are only actually necessary for $\cos(x)$; all other operations could be evaluated in as few as 57 bits and still get the right answer. Moreover, doubling precision on each iteration can overshoot: if we’d started at 64 instead of 80 bits, then we’d try 128, find too-wide an interval, and then have to evaluate every operation at 256 bits. All of this inefficiency comes back to the fact that we are, fundamentally, guessing what precision will be sufficient. It would be better to derive the precision from some analytical method.
That is exactly what Reval does. Suppose we wish to compute $(1-\cos(x))/\sin(x)$ at $x=10^{-12}$, as above, to double precision. Reval first evaluates every operation at a fixed low precision. It then uses the computed intervals to compute amplification factors for each operation. Intuitively, the amplification factor describes how much the operation amplifies error in its inputs. For example, the subtraction operation has an amplification factor of $2^{81}$ because it cancels the first 81 bits of its two inputs. These amplification factors are related to condition numbers and reverse-mode automatic differentiation, though with slight variations to make them sound. The amplification factor (specifically, its logarithm) then indicates how many more bits the inputs of an operator must have for the output to have some target precision. Amplification factors are computed using “exponent tricks”; the subtraction operation’s amplification factor is computed by simply subtracting the unbiased exponents of the high endpoint of $\cos(x)$ and the low endpoint of $1-\cos(x)$. [^3] Plus 1 to account for the significands. Subtracting machine integers, and Section 5 provides similar exponent tricks for other common operations (Section 5).
The amplification factor is then used to assign a precision to each subexpression. For example, the $\sin(x)$ term is assigned a precision of $7878+\log_{2}1=$ 78, while the $\cos(x)$ term is assigned a precision [^4] The details of the computation, including where the values 78 and 83 come from, are discussed in Section 6. of $16483+\log_{2}1+\log_{2}2^{81}=$ 164. This precision assignment is not a guess: it is guaranteed to yield a narrow-enough output interval. So, after one more evaluation with this new precision assignment, Reval computes the correctly-rounded double-precision result.
Deriving a sufficient precision assignment is not always so easy. For example, consider the input $x=10^{-100}$ to the same expression. In its first evaluation at the same fixed precision, Reval computes an interval for the subtraction of $[0,1.033\cdot 10^{25}]$. Because this interval includes 0, Reval cannot bound the exponent of this value, which in turn makes the exponent trick formula return a useless bound of infinity. Reval works around this by guessing $84+512=596$ for the amplification factor (where $512$ is a fixed constant) and ends up assigning 679 bits of precision to the $\cos(x)$ operation. All other operations, however, need no guesses and end up with the starting precision.
Reval now performs a second evaluation at these new precisions. This second evaluation does not yield a narrow-enough interval: Reval’s guess was insufficient. Reval thus re-performs precision assignment. This time the interval for the subtraction is $[4.999\cdot 10^{-101},5.001\cdot 10^{-101}]$ and Reval can compute the amplification factor to be $2^{666}$. Reval then assigns a precision of 749 bits to the $\cos(x)$ term, and re-evaluating again finally yields a correctly-rounded result.
To summarize, Reval initially performs an evaluation at some fixed low precision. It then uses the computed intervals to compute amplification factors, which are then used to compute a precision assignment for each operation. In most cases, the precision assignment guarantees that, after re-evaluation, a narrow-enough interval will be found. [^5] Narrow enough for a “faithful”, though possibly not “correct”, rounding (Brisebarre et al., 2024); see Section 6 However, in other cases, Reval cannot bound the amplification factors and is forced to guess; in this case, it makes guesses with successively larger precisions until a sufficient precision assignment can be computed.
The rest of this paper is organized as follows: Section 3 covers in brief floating-point and interval arithmetic, rounding, and Reval’s mathematical foundations, error Taylor series and condition numbers. Section 4 then derives the amplification factors and shows how they relate to error bounds and precision tuning. Section 5 describes how the amplification factors can be computed quickly, and Section 6 then shows how all of this can be practically implemented.
3. Background
A normal floating-point number is a number ±$\pm(1+m)2^{e}$ where the mantissa $m$ is a value in $[0,1)$ represented with some fixed precision $p$ and where the exponent $e$ is a signed integer. Floating-point numbers also include two zeros (positive and negative), two infinities, a Not-a-Number or NaN value, and a separate class of subnormal values, but in this paper we are largely concerned with normal numbers as other techniques (Flatt and Panchekha, 2023) are available for handling under-flow, over-flow, and the corresponding zero and infinite values. We are specifically concerned with arbitrary-precision floating-point arithmetic as implemented in the well-known MPFR library (Fousse et al., 2007). In MPFR, the precision $p$ can be millions of bits, the exponent is a signed 32-bit integer, and subnormals are not present. [^6] MPFR can emulate subnormals by truncating and shifting, but its internal data representation doesn’t use them.
Rounding
A real number $x$ can be rounded to a floating-point value (at precision $p$). If the real number is already exactly representable, that is easy, but typically $x$ lies strictly between two floating-point values, in which case we say that either one is a faithful rounding of $x$. A faithful rounding of $x$, $\hat{x}$, is typically different from $x$, but at precision $p$ the relative error $(x-\hat{x})/x$ between the two is at most $2^{-p}$, assuming $x$ is within the normal exponent range. A specific rounding mode, such as the default round-to-nearest ties-to-even, may prescribe one of those two values as the correct rounding of $x$. A floating-point implementation $\hat{f}$ of some mathematical function $f$ is said to be faithfully rounded if $\hat{f}(\hat{x})$ is a faithful rounding of $f(\hat{x})$ for all floating-point inputs $\hat{x}$ in its domain, and likewise for correct rounding. Correct rounding is the stricter notion, since most real numbers have two faithful roundings but only one correct rounding. Moreover, correct rounding is deterministic, while faithful rounding is not. For this reason, Reval aims for correctly-rounded results.
A significant advantage of faithful rounding is its relationship with floating-point error. If a number $x$ can be bounded to some range $[x_{1},x_{2}]$, and if the relative error of $x_{1}$ and $x_{2}$ is less than $2^{-p}$, then a faithful rounding of $x$ can be determined. The same is not true of correct rounding: even a very narrow interval $[x_{1},x_{2}]$, much narrower than $2^{-p}$, might still contain the rounding boundary where the correct rounding switches from rounding up to rounding down. For this reason, Reval uses precision tuning to achieve faithful rounding, and then uses an extra “slack” mechanism to handle cases where the result is near a rounding boundary.
Interval Arithmetic
Interval arithmetic performs floating-point computations over intervals instead of single floating-point values. Instead of computing $\hat{f}(x)$, one computes $[y_{\operatorname{lo}},y_{\operatorname{hi}}]=\bar{f}([x_{\operatorname{lo}},x_{\operatorname{hi}}])$ where $y_{\operatorname{lo}}\leq f(x)\leq y_{\operatorname{hi}}$ for all $x_{\operatorname{lo}}\leq x\leq x_{\operatorname{hi}}$. Such an interval is called valid (iee, 2015); the narrowest such interval is called tight. Interval arithmetic is just an abstract domain over floating-point numbers, and the condition $y_{\operatorname{lo}}\leq f(x)\leq y_{\operatorname{hi}}$ is just soundness of the transfer function $\bar{f}$. Many implementations of interval arithmetic exist, both over hardware floating-point numbers (such as Gaol (Goualard, 2019)) and over arbitrary-precision floating-point (such as Rival (Flatt and Panchekha, 2023), used by Reval, or MPFI (Revol and Rouillier, 2005), used by Sollya). These implementations guarantee validity but may depart from tightness in certain difficult cases.
The benefit of interval arithmetic is that intervals are computed using floating-point arithmetic but bound the behavior of real numbers. One valuable application of this property is Ziv’s onion-peeling strategy (Ziv, 1991) for computing the value of a real expression, faithfully-rounded to some target precision $t$. In Ziv’s strategy, the expression is first computed to some interval $[y_{p,\operatorname{lo}},y_{p,\operatorname{lo}}]$ in some working precision $p>t$. The endpoints $y_{p,\operatorname{lo}}$ and $y_{p,\operatorname{hi}}$ are then rounded outward to precision $t$. If the two results are successive floating-point numbers in precision $t$, then either one is a faithful rounding of the expression. Otherwise, the same process can be repeated at higher and higher working precisions $p$.
Error Taylor Series
Error Taylor series bound the difference between a floating-point and real evaluation of the same expression. Consider a sequence of floating-point operations $\hat{z}_{i}=\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})$, where the operations $f_{i}$ may include arithmetic, exponential, trigonometric, or other functions and the computation as a whole represents some expression $\hat{z}$. If the $\hat{f}_{i}$ are faithfully or correctly rounded then we can write $\hat{z}_{i}=f_{i}(\hat{x}_{i},\hat{y}_{i})(1+\varepsilon_{i})$ for some $|\varepsilon_{i}|<2^{-p}$. [^7] More care is needed if the computation of $e$ can be subject to under- or overflow errors. However, this paper does not consider these errors, as prior work (Flatt and Panchekha, 2023) has explained how to handle them in the context of evaluating real expressions. If we perform this rewrite for every operation in the expression, then the value of $\hat{z}$ becomes a real-valued function $\hat{z}(\vec{\varepsilon})$ of many $\varepsilon_{i}$. A real evaluation of $\hat{z}$ then corresponds to $\hat{z}(\vec{0})$, since in a real evaluation there is no rounding error.
Error Taylor series then provide a method to bound the absolute error $\hat{z}(\vec{\varepsilon})-\hat{z}(\vec{0})$. As implemented in tools like FPTaylor (Solovyev et al., 2018) and Satire (Das et al., 2020), this method observes that the $\varepsilon$ values are small and therefore performs a Taylor expansion of $\hat{z}$ in the $\varepsilon$. This expansion is
$\hat{z}(\vec{\varepsilon})=\hat{z}(\vec{0})+\sum_{\varepsilon}\varepsilon\frac{\partial\hat{z}}{\partial\varepsilon}(\vec{0})+O(\varepsilon^{2})$
The first term is the real evaluation; the second term is called the “first-order error”; and the third term is called the “higher-order error”. For example, for the overview example $(1-\cos(x))/\sin(x)$, the first-order error is
$-\frac{\cos(x)}{\sin(x)}\varepsilon_{\cos}+\frac{1-\cos(x)}{\sin(x)}\varepsilon_{-}-\frac{1-\cos(x)}{\sin(x)}\varepsilon_{\sin}+\frac{1-\cos(x)}{\sin(x)}\varepsilon_{/}$
More generally, both the first- and second-order errors can be computed by taking symbolic derivatives of $\hat{z}$. They can then be bounded using a global non-linear optimizer for first-order error and interval arithmetic for second-order error. The two terms use different optimization techniques because the higher-order error is both much smaller—order $\varepsilon^{2}$ instead of $\varepsilon$—and also has quadratically more terms, meaning a faster, looser bound is preferred.
Condition Numbers
Condition numbers simplify the computation of first-order error; while we first learned of the connection in ATOMU (Zou et al., 2019), the basic insight is the same as that in reverse-mode automatic differentiation. Suppose one wishes to bound the relative error of a floating-point vs real computation of an expression $\hat{z}$. Then, following the error Taylor series idea above, one needs to compute the first-order error divided by $\hat{z}$, which involves computing
$\frac{(\partial\hat{z}/\partial\varepsilon_{i})(\vec{0})}{\hat{z}(\vec{0})}.$
Condition numbers rearrange this computation into a sum of products
$\sum_{p}\prod_{f_{j}\in p}\Gamma_{f_{j}}.$
Here, the sum ranges over all paths $p$, where a path means a path in the expression tree from the root to some intermediate value. For example, in the overview example $(1-\cos(x))/\sin(x)$, the path from the root to $\cos(x)$ involves a division and a subtraction, while the path from the root to $\sin(x)$, involves just a division. Then the $f_{j}$ ranges over each operation in that path, and for each path their condition numbers $\Gamma_{f_{j}}$ are multiplied together. The condition number $\Gamma_{f}$ is equal to $xf^{\prime}(x)/f(x)$, which can be evaluated symbolically for all operations $f_{i}$ in the expression $\hat{z}$. The benefit of rearranging error Taylor series in this way is that the $\Gamma_{f_{j}}$ operations are local to specific steps in the expression $\hat{z}$, while derivatives $\partial\hat{z}/\partial\varepsilon_{i}$ involve the whole expression $\hat{z}$ at once. That can allow, as in ATOMU (Zou et al., 2019), identifying specific operations that result in high error.
4. Fast Sound Precision Tuning
Reval’s precision assignment algorithm is based on error bounds for floating-point computations. At a high level, these error bounds are based on error Taylor series and condition numbers, with some variations to ensure soundness and performance. Their derivation is fairly technical; we invite readers uninterested in the technical details to focus on the high-level description in Section 4.1 and then on Equations (1), (2), and (3), which are the key results proved in the rest of this section.
4.1. High-level Approach
Intuitively, the endpoints computed in interval arithmetic bound all possible faithfully-rounded evaluations of a given expression. Sound bounds on faithfully-rounded evaluations thus indicate how wide that interval may be. If that error bound is parameterized by the precision of each operation, then the sound error bound is an inequality that we can solve for an appropriate precision assignment. Importantly, these error bounds need to be sound, always over-approximating the error; otherwise, a too-low precision will be assigned and the resulting output interval will be too wide. But if the error is over-approximated by too much, too-high a precision will be assigned, which will make evaluation take too long.
The need for soundness and accuracy suggests error Taylor series, but performance is a problem. Error Taylor series are only sound if the “higher-order error” term is included, and that term is expensive to compute: it requires computing quadratically-many second derivatives of the expression. We instead introduce a novel variation on error Taylor series that achieves soundness using only first derivatives. Equation 1 shows that variation. Even this variation, computed naively, would take $O(n^{2})$ time to compute (where $n$ is the number of real operations), but an algebraic reorganization inspired by (and related to) condition numbers allows computing it in $O(n)$ time. That reorganization is shown in Equation 2. Finally, the error bound can be solved for an appropriate precision assignment; the solution is shown in Equation 3.
4.2. Sound First-Order Error
We first derive sound floating-point error bounds that only use first, not second, derivatives. The bound is derived similarly to standard error Taylor series, but applies the Lagrange remainder theorem differently to achieve soundness without second derivatives.
Analogously to error Taylor series, consider a fixed sequence $\hat{z}_{i}=\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})$, of floating-point operations $\hat{f}_{i}$ on input and output floating-point registers $\hat{x}_{i}$, $\hat{y}_{i}$, and $\hat{z}_{i}$; we assume each $\hat{x}_{i}$ is itself the output register $\hat{z}_{j}$ of some prior operation. Naturally, some operations have one or even zero arguments (for floating-point constants like $\pi$), but for simplicity we write $\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})$ for each operation. Each operation also has a corresponding real-number operation $z_{i}=f(x_{i},y_{i})$. We also assume each $\hat{f}_{i}$ is faithfully rounded,888 Reval’s interval arithmetic is not tight, especially in challenging cases like $\cos(\pi)$, but adding a couple of extra bits allows for a couple of ULPs of looseness; see Section 6 for more. so that
$\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})=f_{i}(\hat{x}_{i},\hat{y}_{i})(1+\varepsilon_{i}),$
for some $|\varepsilon_{i}|\leq 2^{-p_{i}}$, where $p_{i}$ is the precision used for this operation.999 This formula is not correct in the presence of under- and over-flow, though many cases of such are handled using movability flags (Flatt and Panchekha, 2023), a feature of the interval library we use. The remaining cases are handled by the slack mechanism of Section 6. Substituting in $\hat{x}_{i}$ and $\hat{y}_{i}$, which are themselves computed in prior instructions, yields $\hat{z}_{i}$ as a function of all of the $\varepsilon_{j}$s, which we now write $\hat{z}_{i}(\vec{\varepsilon})$.
The relative error of $\hat{z}_{i}$, then, is a function of the $\vec{\varepsilon}$. Error Taylor series provide a way to bound this relative error. Notice that $\hat{z}_{i}(\vec{0})=z_{i}$; the numerator of the relative error formula is thus $\hat{z}_{i}(\vec{\varepsilon})-\hat{z}_{i}(\vec{0})$. But $\hat{z}_{i}(\vec{0})$ is the first term of the Taylor series for $\hat{z}_{i}$.
Here we diverge from error Taylor series, applying the Lagrange remainder theorem at order 1 instead of order 2: [^10] There are many variations of the Lagrange remainder theorem; the statement below is one of them but can also be seen as a corollary of the mean value theorem.
$\hat{z}_{i}(\vec{\varepsilon})-\hat{z}_{i}(\vec{0})=\sum_{j\leq i}\varepsilon_{j}\frac{\partial\hat{z}_{i}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})$
for some $0\leq\vec{\varepsilon^{*}}\leq\vec{\varepsilon}\leq 2^{-p_{i}}$ (or the symmetrical bound for $\varepsilon_{j}<0$). In this bound, $\vec{\varepsilon}$ represents the rounding error of a faithfully-rounded execution, so $\vec{\varepsilon^{*}}$ represents the rounding error of some other, also faithfully-rounded execution (of $\partial f/\partial\varepsilon_{j}$).
We don’t know the signs of the $\varepsilon_{i}$, just a bound on their magnitude, but using the triangle inequality we can still bound the relative error of $\hat{z}_{i}$:
(1) $\left|\frac{\hat{z}_{i}-z_{i}}{z_{i}}\right|=\left|\frac{\hat{z}_{i}(\vec{\varepsilon})-\hat{z}_{i}(\vec{0})}{z_{i}}\right|\leq\max_{\vec{\varepsilon^{*}}}\sum_{j\leq i}2^{-{p_{j}}}\left|\frac{1}{z_{i}}\frac{\partial\hat{z}_{i}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|.$
This is similar to the first-order error in standard error Taylor series, except that Equation 1 maximizes the error over all $\vec{\varepsilon^{*}}$, where standard first-order error just considers $\vec{\varepsilon^{*}}=\vec{0}$. This is what ensures soundness, replacing the higher-order error term used in standard error Taylor series. At the core, this difference comes from error Taylor series applying the Lagrange remainder theorem to the second-order Taylor term, instead of the first-order Taylor term as done here.
4.3. Linear-Time Error Bounds
Next, we consider the algorithmic complexity of computing Equation 1. This equation involves a sum of $O(i)$ terms. Moreover, each term requires computing a different derivative of $\hat{z}_{i}$. Since $\hat{z}_{i}$ is an expression containing (at most) $i$ operations, that derivative can be computed in $O(i)$ time using forward-mode automatic differentiation (Das et al., 2020). However, that still results in an overall runtime of $O(i^{2})$. This is unacceptable for Reval, since actually executing the instructions takes only $O(i)$ time. This issue can be addressed by rearranging the formula algebraically; the derivation is similar to condition numbers or reverse-mode automatic differentiation.
Consider the $\partial\hat{z}_{i}/\partial\varepsilon_{j}$ term and substitute in $\hat{z}_{i}=f_{i}(\hat{x}_{i},\hat{y}_{i})(1+\varepsilon_{i})$. Then we have two cases: either $i=j$ or $i>j$. If $i=j$, we have
$\frac{\partial\hat{z}_{i}}{\partial\varepsilon_{i}}=f_{i}(\hat{x}_{i},\hat{y}_{i}),$
because the $\hat{x}_{i}$ and $\hat{y}_{i}$ terms were computed in previous operations and are thus independent of $\varepsilon_{i}$. On the other hand, if $i>j$, then $\varepsilon_{i}$ is independent of $\varepsilon_{j}$, so we have
$\frac{\partial\hat{z}_{i}}{\partial\varepsilon_{j}}=(1+\varepsilon_{i})\left((\partial_{1}f)(\hat{x}_{i},\hat{y}_{i})\frac{\partial\hat{x}_{i}}{\partial\varepsilon_{j}}+(\partial_{2}f)(\hat{x}_{i},\hat{y}_{i})\frac{\partial\hat{y}_{i}}{\partial\varepsilon_{j}}\right),$
where $\partial_{1}f$ and $\partial_{2}f$ refer to the derivatives of $f$ relative to its first and second arguments. Performing some rearrangements and noting that $|\varepsilon_{i}|<2^{-p_{i}}$, we now have:
$\frac{\hat{z}_{i}-z_{i}}{z_{i}}\leq\overbrace{\max_{\vec{\varepsilon^{*}}}\sum_{j\leq i}2^{-p_{j}}\left|\frac{1}{z_{i}}\frac{\partial\hat{z}_{i}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|}^{\operatorname{bound}(\hat{z}_{i})}$ $\leq 2^{-p_{i}}\overbrace{\max_{\vec{\varepsilon^{*}}}\left|\frac{f_{i}(x^{*}_{i},y^{*}_{i})}{f_{i}(x_{i},y_{i})}\right|}^{\operatorname{intro}(\hat{z}_{i})}$
$+(1+2^{-p_{i}})\left(\max_{\vec{\varepsilon^{*}}}\left|\frac{x_{i}(\partial_{1}f_{i})(x^{*}_{i},y^{*}_{i})}{f_{i}(x_{i},y_{i})}\right|\right)\left(\max_{\vec{\varepsilon^{*}}}\sum_{j<i}2^{-p_{j}}\left|\frac{1}{x_{i}}\frac{\partial\hat{x_{i}}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|\right)$
$+(1+2^{-p_{i}})\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\left|\frac{y_{i}(\partial_{2}f_{i})(x^{*}_{i},y^{*}_{i})}{f_{i}(x_{i},y_{i})}\right|\right)}_{\operatorname{ampl}_{2}(\hat{z}_{i})}\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\sum_{j<i}2^{-p_{j}}\left|\frac{1}{y_{i}}\frac{\partial\hat{y_{i}}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|\right)}_{\operatorname{bound}(\hat{y}_{i})}$
where $x^{*}_{i}$ and $y^{*}_{i}$ are $\hat{x}_{i}$ and $\hat{y}_{i}$ evaluated at $\vec{\varepsilon^{*}}$. Rewriting this gargantuan formula using the helper functions defined by the braces, we get:
(2) $\operatorname{bound}(\hat{z})\leq 2^{-p}\operatorname{intro}(\hat{z})+(1+2^{-p})\operatorname{ampl}_{1}(\hat{z})\operatorname{bound}(\hat{x})+(1+2^{-p})\operatorname{ampl}_{2}(\hat{z})\operatorname{bound}(\hat{y}),$
where all variables have the same subscript, $i$. Equation 2 can be evaluated in a single linear pass through the program, which gives it the required linear algorithmic complexity, assuming the $\operatorname{intro}$ and $\operatorname{ampl}_{k}$ terms can be evaluated in $O(1)$ time.
The intuition behind this formula is that every operation $\hat{z}_{i}$ introduces some error $\operatorname{intro}(\hat{z}_{i})$ and also amplifies error in its inputs by factors of $\operatorname{ampl}_{k}(\hat{z}_{i})$. The error bound for $\hat{z}_{i}$ just sums these three factors. Readers familiar with condition numbers will immediately see the similarity between Equation 2 and the condition number formula $e_{z}=\mu+\Gamma_{x}e_{x}+\Gamma_{y}e_{y}$ (Zou et al., 2019). However, condition numbers bound only the first-order error so are unsound, while Equation 2 is sound. This in turn causes differences like the $\operatorname{intro}(\hat{z}_{i})$ term in place of $\mu$, the slight difference in definition between $\operatorname{ampl}_{k}(\hat{z}_{i})$ and $\Gamma_{k}$, and the extra $1+2^{-p_{i}}$ term. These differences are small but necessary for soundness. [^11] In fact, the name $\operatorname{ampl}$ is a pun on both “amplification factor” and ”ample” as in sound.
4.4. Precision Tuning
Next, we treat Equation 2 as an inequality that can be solved for the $p_{j}$ terms. That is, we seek $p_{j}$, for all $j$, such that $\operatorname{bound}(\hat{z}_{i})\leq 2^{-t}$ for some target precision $t$. To do so, split the allowed $2^{-t}$ error equally among the three terms on the right hand side of Equation 2:
$\operatorname{bound}(\hat{z}_{i})\leq 2^{-t}$ $\Leftarrow 2^{-p_{i}}\operatorname{intro}(\hat{z}_{i})\leq\frac{1}{3}2^{-t}$
$\land(1+2^{-p_{i}})\operatorname{ampl}_{1}(\hat{z}_{i})\operatorname{bound}(\hat{x}_{i})\leq\frac{1}{3}2^{-t}$
$\land(1+2^{-p_{i}})\operatorname{ampl}_{2}(\hat{z}_{i})\operatorname{bound}(\hat{y}_{i})\leq\frac{1}{3}2^{-t},$
where the left arrow is a reversed logical implication.
Let’s take the first conjunction, bounding the introduced error, first. Rearranging a little and noting that $\log_{2}3\leq 2$, we have $p_{i}\geq t+2+\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil$. Likewise consider, say, the first amplified error term. It can be rearranged into $-\log_{2}\operatorname{bound}(\hat{x}_{i})\geq t+\log_{2}(3+3\cdot 2^{-p_{i}})+\log_{2}\operatorname{ampl}_{1}(\hat{z}_{i})$. As long as $p_{i}\geq 2$, the term $\log_{2}(3+3\cdot 2^{-p_{i}})\leq 2$ so we must bound $\hat{x}_{i}$’s error to at most $t+2+\log_{2}\operatorname{ampl}_{1}(\hat{z}_{i})$ bits.
Combining these, we have:
(3) $-\log_{2}\operatorname{bound}(\hat{z}_{i})\geq t\Longleftarrow\bigwedge\left[\begin{array}[]{l}p_{i}=\max(2,t+2+\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil)\\ -\log_{2}\operatorname{bound}(\hat{x}_{i})\geq t+2+\lceil\log_{2}\operatorname{ampl}_{1}(\hat{z}_{i})\rceil\\ -\log_{2}\operatorname{bound}(\hat{y}_{i})\geq t+2+\lceil\log_{2}\operatorname{ampl}_{2}(\hat{z}_{i})\rceil\end{array}\right.$
Intuitively, the root operation must be evaluated at precision $t$ plus a few extra bits, and then each subexpression’s target precision must be higher to accommodate the amplification of its error.
5. The Exponent Trick
To actually use Equation 2 we must compute $\operatorname{intro}(\hat{z}_{i})$ and $\operatorname{ampl}_{k}(\hat{z}_{i})$, which requires optimizing over $\varepsilon^{*}$. However, we can use interval arithmetic to perform this optimization. Moreover, already-computed interval outputs from each operation can be reused to make this fast.
5.1. Optimizing Over $\vec{\varepsilon^{*}}$
Interval arithmetic can perform the optimization over $\vec{\varepsilon^{*}}$. Consider $\operatorname{intro}(\hat{z}_{i})$ and suppose we already have an interval evaluation, $\bar{z}_{i}$, of $z_{i}$. The very nature of interval arithmetic is that the interval bounds all faithfully-rounded evaluations of $z_{i}$, [^12] At equal or higher precision, an invariant Reval maintains (Section 6). meaning $\bar{z}_{i}$ contains both $z_{i}$ and $z_{i}^{*}$. So,
$\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{f(x_{i}^{*},y_{i}^{*})}{f(x_{i},y_{i})}\right|\right\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{z_{i}^{*}}{z_{i}}\right|\right\rceil\in\lceil\log_{2}|\bar{z}_{i}/\bar{z}_{i}|\rceil$
Note that $\bar{z}_{i}/\bar{z}_{i}$ is an interval-arithmetic division, so it isn’t equal to 1; it’s an interval that contains 1, but also some larger and smaller values. Likewise, the absolute value, logarithm, and ceiling operators are all interval operators. In other words, $\operatorname{intro}$—and, with a similar trick, $\operatorname{ampl}_{k}$—can be computed with standard interval arithmetic.
The upper bound on $\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil$ is the high endpoint of $\lceil\log_{2}|\bar{z}_{i}/\bar{z}_{i}|\rceil$. Write $\operatorname{lo}(\bar{x})$ and $\operatorname{hi}(\bar{x})$ for the high and low endpoint of an interval $\bar{x}$, and assume for now (see Section 6) that $\bar{x}$ contains neither zero nor infinity. Then,
$\operatorname{hi}(\lceil\log_{2}|\bar{z}_{i}/\bar{z}_{i}|\rceil)\leq\lceil\log_{2}\operatorname{hi}(|\bar{z}_{i}|)\rceil-\lfloor\log_{2}\operatorname{lo}(|\bar{z}_{i}|)\rfloor\leq\underbrace{\lfloor\log_{2}\operatorname{hi}(|\bar{z}_{i}|)\rfloor+1}_{\operatorname{maxlog}\bar{z}_{i}}-\underbrace{\lfloor\log_{2}\operatorname{lo}(|\bar{z}_{i}|)\rfloor.}_{\operatorname{minlog}\bar{z}_{i}}$
The term $\operatorname{maxlog}\bar{z}_{i}-\operatorname{minlog}\bar{z}_{i}$ comes up repeatedly, so we shorten it to $\operatorname{logspan}\bar{z}_{i}$.
Importantly, $\lfloor\log_{2}x\rfloor$, for a floating-point number $x$, is just the exponent in the floating-point representation of $x$, which in MPFR is stored as a signed machine integer. [^13] Technically, in MPFR, the exponent is offset by 1 from this definition, which Reval corrects for. And $\bar{z}_{i}$ is simply the already-computed interval output of the operation $z_{i}=f(x_{i},y_{i})$, so its exponent can be accessed quickly. In other words, $\operatorname{minlog}$ and $\operatorname{maxlog}$ can be computed extremely quickly, requiring no library calls, pointer chasing, or allocation.
We call this the “exponent trick”: bounding the logarithms of $\operatorname{intro}$ and $\operatorname{ampl}_{k}$ using only: 1) integer operations, 2) on $\operatorname{minlog}$ and $\operatorname{maxlog}$, 3) of already-computed intervals $\bar{x}$, $\bar{y}$, and $\bar{z}$. The remainder of this section describes exponent tricks for various operators. Some of the derivations are mathematically challenging, but all the key results are shown in Tables 2 and 2.
5.2. Simple amplification factors
Just like $\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil$, $\lceil\log_{2}\operatorname{ampl}_{k}(\hat{z}_{i})\rceil$ can be computed using exponent tricks. However, there is an additional challenge, because the definition of $\operatorname{ampl}_{k}$ refers to $(\partial_{k}f_{i})(x_{i}^{*},y_{i}^{*})$. Its value is not known, and the interval $(\partial_{k}f_{i})(\bar{x}_{i},\bar{y}_{i})$ that contains it has typically not been computed. Different operators use different techniques to overcome this challenge.
For addition and subtraction, there’s no problem at all: $\log_{2}\operatorname{ampl}_{1}(\hat{x}_{i}+\hat{y}_{i})$ is just equal to $\log_{2}|x_{i}/z_{i}|$, which is bounded by $\operatorname{maxlog}\bar{x}_{i}-\operatorname{minlog}\bar{z}_{i}$. Subtraction likewise presents no difficulty.
For multiplication, an extra $\operatorname{logspan}$ factor appears: $\log_{2}\operatorname{ampl}_{1}(\hat{x}_{i}\cdot\hat{y}_{i})$ is equal to $\log_{2}|x_{i}y_{i}^{*}/x_{i}y_{i}|$; the $x_{i}$ terms cancel leaving a bound of $\operatorname{logspan}\bar{y}_{i}$.
Similarly, the amplification factor for the power function in its first argument,
$\operatorname{ampl}_{1}(\operatorname{pow}(\hat{x}_{i},\hat{y}_{i}))=\left|\frac{x_{i}y_{i}^{*}\operatorname{pow}(x_{i}^{*},y_{i}^{*}-1)}{\operatorname{pow}(x_{i},y_{i})}\right|=\left|\frac{x_{i}y_{i}^{*}\operatorname{pow}(x_{i}^{*},y_{i}^{*})}{x_{i}^{*}\operatorname{pow}(x_{i},y_{i})}\right|=\left|y_{i}^{*}\frac{x_{i}}{x_{i}^{*}}\frac{z_{i}^{*}}{z_{i}}\right|,$
results in the bound $\operatorname{maxlog}\bar{y}_{i}+\operatorname{logspan}\bar{x}_{i}+\operatorname{logspan}\bar{z}_{i}$.
Division, square and cube roots, exponents, and logarithms allow a similar algebraic rearrangement, producing a formula that uses only integer operations on $\operatorname{maxlog}$ and $\operatorname{minlog}$ of $\bar{x}_{i}$, $\bar{y}_{i}$, and $\bar{z}_{i}$. The resulting formulas are shown in Table 2.
5.3. Approximations
However, for some operators, an approximation has to be performed to bound $(\partial_{k}f)(x,y)$ via some expression over $x$, $y$, or $z$. This approximation allows us to construct an exponent tricks formula for $\operatorname{ampl}_{k}(f)$, that is, a formula that bounds $\lceil\log_{2}\operatorname{ampl}_{k}(f)\rceil$ using only integer operations on $\operatorname{maxlog}$ and $\operatorname{minlog}$ of $\bar{x}$, $\bar{y}$, and $\bar{z}$. Table 2 shows all of the approximations used and the exponent trick formulas they result in, and readers are invited to focus their attention on it. The rest of this section merely contains detailed derivations for each row of that table.
Sine and Cosine
The amplification factor for $\sin(x)$ involves $\cos(x)$. Because $|\cos(x)|\leq 1$, we have $\log_{2}|\cos(x^{*})|\leq 0$; this allows us to show that
$\lceil\log_{2}\operatorname{ampl}_{1}(\sin(\hat{x}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{y}.$
Likewise, for cosine we must bound $\sin(x)$; one simple approximation is $|\sin(x)|\leq\min(|x|,1)$, meaning $\log_{2}|\sin(x^{*})|\leq\min(\operatorname{maxlog}(\bar{x}),0)$. This produces the approximation
$\lceil\log_{2}\operatorname{ampl}_{1}(\cos(\hat{z}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{y}+\min(\operatorname{maxlog}(\bar{x}),0).$
These approximations are sound and easy to compute. Moreover, the approximation error is small when either the input or the output is near zero, which helps reduce the needed precision.
Power
The power function $x^{y}$ has an amplification factor for $y$ that uses a $\log(x^{*})$ term. This means that we must bound $\log_{2}|\log(x)|$ in terms of $\log_{2}|x|$. To do so, we first rewrite
$\log_{2}|\log(x)|=\log_{2}\left|\frac{\log_{2}(x)}{\log_{2}(e)}\right|=\log_{2}|\log_{2}(x)|-\log_{2}\log_{2}e\leq\log_{2}|\log_{2}(x)|-\frac{1}{2}.$
Set $s=\log_{2}(x)$; since $\log_{2}|s|\leq|s|-\frac{1}{2}$, we have $\log_{2}|s|-\frac{1}{2}\leq|s|-1$. This ultimately means that $\log_{2}|\log(x^{*})|$ is bounded above by $|\log_{2}|\bar{x}||-1$. That is just $\max(|\operatorname{maxlog}(\bar{x})|,|\operatorname{minlog}(\bar{x})|)-1$.
Hyperbolic Functions
The amplification factors for $\sinh$ and $\cosh$ both require bounding $\tanh(x)$. In both cases, we use the well-known fact that $|\tanh(x)|\leq\min(|x|,1)$, leading to $\lceil\log_{2}|\tanh(x)|\rceil\leq\min(\operatorname{maxlog}(\bar{x}),0)$. This approximation is relatively tight, producing at most one bit of error.
The amplification factor for $\tanh(x)$ requires bounding $x/\sinh(x)\cosh(x)$. We use a very weak bound, $x/\sinh(x)\cosh(x)\leq 1$. This over-approximates significantly for large values of $x$, though not near 0.
Tangent
The amplification factor for $\tan(\hat{x}_{i})$ is $|x_{i}/\tan(x_{i})\cos^{2}(x_{i}^{*})|$. To bound its logarithm, we bound $\cos(x)$ in terms of $\tan(x)$. Since $\sin(x)^{2}+\cos(x)^{2}=1$, $\sqrt{1/2}\leq\max(|\sin(x)|,|\cos(x)|)\leq 1$. Divide both sides by $|\cos(x)|$ and take the logarithm; this yields:
$-\frac{1}{2}-\log_{2}|\cos(x)|\leq\max(\log_{2}|\tan(x)|,0)\leq-\log_{2}|\cos(x)|,$
or, in other words,
$-2\log_{2}|\cos(x)|\in 2\max(\log_{2}(|\tan(x)|),0)+[0,1],$
leaving
$\lceil\log_{2}\operatorname{ampl}_{1}(\tan(\hat{x}_{i}))\rceil$ $\leq\operatorname{maxlog}\bar{x}_{i}-\operatorname{minlog}\bar{z}_{i}+2\max(\operatorname{maxlog}\bar{z}_{i},0)+1$
$=\operatorname{maxlog}\bar{x}_{i}-\operatorname{minlog}\bar{z}_{i}+\operatorname{maxlog}\bar{z}_{i}+\max(\operatorname{maxlog}\bar{z}_{i},-\operatorname{maxlog}\bar{z}_{i})+1$
$=\operatorname{maxlog}\bar{x}_{i}+|\operatorname{maxlog}\bar{z}_{i}|+\operatorname{logspan}\bar{z}_{i}+1,$
which now only uses already-computed intervals. Moreover, this approximation is off by at most one bit.
Arctangent
The amplification factor for the arctangent contains a $x/(x^{2}+1)$ term, which we must bound in terms of $x$. Some rearrangement yields
$\log_{2}\left|\frac{x}{x^{2}+1}\right|=-\log_{2}\left|x+\frac{1}{x}\right|.$
Now consider $x\geq 1$; for these inputs, we clearly have
$\log_{2}|x+1/x|\geq\log_{2}|x|=\left|\log_{2}|x|\right|$
Since the left- and right-hand sides are both invariant under $x\mapsto 1/x$ and $x\mapsto-x$, this inequality is then true for all $x$. Moreover, this approximation is off by at most one bit.
Operation ($f$) Amplification factor ($\operatorname{ampl}_{k}(f)$) Exponent tricks bound for $\lceil\log_{2}\operatorname{ampl}_{k}\rceil$
$x+y$ $k=1:|x/(x+y)|$ $k=2:|y/(x+y)|$ $k=1:\operatorname{maxlog}(\bar{x})-\operatorname{minlog}(\bar{z})$ $k=2:\operatorname{maxlog}(\bar{y})-\operatorname{minlog}(\bar{z})$
$x-y$ $k=1:|x/(x+y)|$ $k=2:|y/(x+y)|$ $k=1:\operatorname{maxlog}(\bar{x})-\operatorname{minlog}(\bar{z})$ $k=2:\operatorname{maxlog}(\bar{y})-\operatorname{minlog}(\bar{z})$
$x\times y$ $k=1:|y^{*}/y|$ $k=2:|x^{*}/x|$ $k=1:\operatorname{logspan}(\bar{y})$ $k=2:\operatorname{logspan}(\bar{x})$
$x\div y$ $k=1:|y/y^{*}|$ $k=2:\left|\frac{yx}{{y^{*}}^{2}}/\frac{x}{y}\right|$ $k=1:\operatorname{logspan}(\bar{y})$ $k=2:\operatorname{logspan}(\bar{x})+2(\operatorname{logspan}(\bar{y}))$
$\sqrt{x}$ $|1/2\sqrt{x/x^{*}}|$ $\frac{1}{2}\operatorname{logspan}(\bar{x})-1$
$3]{x}$ $|1/3(x/x^{*})^{2/3}|$ $\frac{2}{3}\operatorname{logspan}(\bar{x})-1$
$\log(x)$ $\left|(x/x^{*})/\log(x)\right|$ $\operatorname{logspan}(\bar{x})-\operatorname{minlog}(\bar{z})$
$e^{x}$ $|xe^{x^{*}}/e^{x}|$ $\operatorname{maxlog}(\bar{x})+\operatorname{logspan}(\bar{z})$
$x^{y}$, $k=1$ $\left|xy^{*}(x^{*})^{y^{*}-1}/x^{y}\right|$ $\operatorname{maxlog}(\bar{y})+\operatorname{logspan}(\bar{x})+\operatorname{logspan}(\bar{z})$
Table 1. Bounds on amplification factors using only the exponents of already-computed intervals $\bar{x}$, $\bar{y}$, and $\bar{z}$. The formulas in this table do not use any approximations.
Operator ($f$) Approximation used & exponent tricks bound
$x^{y}$, $k=2$ $\log_{2}|x|\leq|x|-\frac{1}{2}$
$\operatorname{maxlog}(\bar{y})+\max(|\operatorname{minlog}(\bar{x})|,|\operatorname{maxlog}(\bar{x})|)-1+\operatorname{logspan}(\bar{z})$
$\cos(x)$ $|\sin(x)|\leq\min(|x|,1)$
$\operatorname{maxlog}(\bar{x})-\operatorname{minlog}(\bar{z})+\min(\operatorname{maxlog}(\bar{x}),0)$
$\sin(x)$ $|\cos(x)|\leq 1$
$\operatorname{maxlog}(\bar{x})-\operatorname{minlog}(\bar{z})$
$\tan(x)$ $\max(|\sin(x)|,|\cos(x)|)\geq\frac{1}{2}$
$\operatorname{maxlog}(\bar{x})+\max(|\operatorname{minlog}(\bar{z})|,|\operatorname{maxlog}(\bar{z})|)+\operatorname{logspan}(\bar{z})+1$
$\cosh(x)$ $|\tanh(x)|\leq\min(|x|,1)$
$\operatorname{maxlog}(\bar{x})+\operatorname{logspan}(\bar{z})+\min(\operatorname{maxlog}(\bar{x}),0)$
$\sinh(x)$ $|\tanh(x)|\leq\min(|x|,1)$
$\operatorname{maxlog}(\bar{x})+\operatorname{logspan}(\bar{z})-\min(\operatorname{minlog}(\bar{x}),0)$
$\tanh(x)$ $|x/\sinh(x)\cosh(x)|\leq 1$
$\operatorname{logspan}(\bar{z})+\operatorname{logspan}(\bar{x})$
$\operatorname{atan}(x)$ $\log_{2}(|x+\frac{1}{x}|)\geq|\log_{2}(|x|)|$
$\operatorname{logspan}(\bar{x})-\min(|\operatorname{minlog}(\hat{x})|,|\operatorname{maxlog}(\hat{x})|)-\operatorname{minlog}(\hat{z})$
[…]
$\operatorname{maxlog}(\bar{x})+\operatorname{maxlog}(\bar{y})-2\max(\operatorname{minlog}(\bar{x}),\operatorname{minlog}(\bar{y}))-\operatorname{minlog}(\bar{z})$
Table 2. Exponent trick bounds derived via approximation. Most approximations, including for $\tan$, $\operatorname{atan}$, and $\cosh$, over-approximate by at most a factor of two. The bounds for $\sin$ and $\cos$ over-approximate by more, though on relatively rare inputs.
6. Implementation Details
Reval provides a simple user interface: the user first compiles an expression, which may have free variables, and then applies it to specific values. The compilation produces a simple register machine with a vector of instructions and separate vectors with, for each instruction, an output interval and a precision assignment.
6.1. Compilation
Reval supports expressions with real-number operations, and real constants, boolean operations, and if statements, but no loops. Specifically, Reval supports the constants $\pi$ and $e$, the standard arithmetic operators (plus, minus, times, divide), and every function specified in the POSIX standard for math.h, including square and cube roots; exponents, powers, and logarithms; trigonometric, arc-trigonometric, and hyperbolic and arc-hyperbolic functions; the error and gamma functions; modulus and remainder operators; rounding functions; and miscellaneous others like $\min$ and $\max$. It also supports comparison functions and boolean connectives. Thanks to the lack of loops, each expression can be flattened to a linear instruction sequence. Users can also supply a batch of multiple expressions to evaluate on the same inputs. Reval performs common subexpression elimination so this can dramatically reduce runtime if the expressions have shared structure, though this feature is not used in this paper.
Instructions are written in two-argument form, reading in their arguments from input registers and returning a result that is saved in the instruction’s output register. Interval operators come from the Rival library (Flatt and Panchekha, 2023), which has a couple of unusual features. It provides boolean functions like comparisons and connectives that output ternary “boolean intervals”; for example, $[1,1]<[0,2]$ returns the “maybe” interval [,]bottomtop[$\bot$,$\top$][ ⊥ , ⊤ ]. An if statement with a “maybe” condition unions the intervals in both branches. In the presence of domain errors, Rival intervals enclose all outputs on valid inputs; for example, $\sqrt{[-1,1]}=[0,1]$. A special empty interval $\emptyset$ is used when the whole input interval is invalid. All intervals also track domain errors using a ternary error flag; for $\sqrt{[-1,-1]}$ the error flag is $\top$ and for $\sqrt{[-1,1]}$ the error flag is [,]bottomtop[$\bot$,$\top$][ ⊥ , ⊤ ]. Rival also provides movability flags (Flatt and Panchekha, 2023) that detect some unevaluable inputs.
Reval typically attempts to correctly round each expression to double precision. However, this rounding behavior is customizable by supplying a discretization. A discretization is a type $T$, a target precision $t$, a conversion function from MPFR values to $T$s, and a distance function on pairs of $T$s. The distance function is called with the two converted $T$ endpoints of an interval and returns one of three values: the endpoints can be equal, unequal, or neighbors, meaning that the two $T$ values share a rounding boundary. The “equal” return value is used for correct rounding; if the user instead wants faithful rounding, they can supply a distance function that returns “neighbor” instead of “equal”.
6.2. Evaluation
Reval evaluates each instruction in turn, at that instruction’s precision assignment. Then, for each output register, Reval uses the conversion and distance functions to determine if the computed intervals are sufficiently narrow. If not, Reval performs a precision assignment using the exponent trick formulas and the already-computed intervals for each register, and repeats this process until a sufficiently-narrow interval is found. This minimal evaluation loop reduces evaluation overhead, and performance is largely determined by the choice of precision for each instruction.
If the precision assigned to a specific instruction does not change (or even decreases) from one iteration to the next, there’s no need to recompute that instruction, as long as its inputs are not recomputed. Reval checks for this condition during precision assignment and sets a bit in a per-instruction “repeat vector”; instructions with the repeat bit set are not re-executed. On the other hand, if an operation is ever assigned a precision higher than a user-specified precision bound, Reval terminates execution and signals an error. This “early stopping” can trigger as early as the first iteration and saves significant time on points Reval cannot evaluate. Reval can also exit the evaluation loop if the output intervals’ error flag is set to $\top$, if movability flags prove the expression unevaluable, or if a user-defined iteration limit is reached.
6.3. Precision Tuning
Reval’s precision tuning pass is basically Equation 3 converted to code. Reval allocates a vector $P[z_{i}]$ and fills the vector starting from the end. The already-filled, later elements of the array represent the precisions $p_{i}$ for each instruction, while the initial, un-filled elements represent a target precision for $\operatorname{bound}(\hat{z}_{i})$. In code, the algorithm is:
$P[z_{n}]=t$
[…]
$\quad P[z_{i}]:=\max(2,P[z_{i}]+2+\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil)$
$\quad P[x_{i}]:=\max(P[x_{i}],P[z_{i}]+2+\lceil\log_{2}\operatorname{ampl}_{1}(\hat{z}_{i})\rceil)$
$\quad P[y_{i}]:=\max(P[y_{i}],P[z_{i}]+2+\lceil\log_{2}\operatorname{ampl}_{2}(\hat{z}_{i})\rceil)$
[…]
The $\operatorname{intro}$ and $\operatorname{ampl}$ terms are computed using the exponent trick formulas of Tables 2 and 2. This single “backward pass” over the precision array suffices to compute a precision assignment for each instruction. Reval performs some additional, minor optimizations to make this formula even faster to compute, including pre-propagating the “${}+2$” factors and computing the $\operatorname{intro}$ and $\operatorname{ampl}$ terms simultaneously.
Soundness ensures that Reval terminates: without a sound precision assignment, it would be possible to enter an infinite loop, where at each iteration the same precisions are assigned. Reval does include code to check for this condition (in case the user provides a custom discretization with a too-low target precision $t$) but this code is not triggered in our evaluation. However, achieving termination also requires some additional pragmatic changes to the simple theory of Equation 3.
6.4. Slack
Sometimes, an interval crosses zero, in which case $\operatorname{minlog}(\bar{z})$ is not defined—it approaches negative infinity. Similarly, some intervals are unbounded, in which case $\operatorname{maxlog}(\bar{z})$ approaches positive infinity. In both cases, the exponent trick formulas would evaluate $\operatorname{intro}$ and $\operatorname{ampl}_{k}$ as infinity, which of course is not a usable precision assignment. Intuitively, what’s happening is that the precision assignment algorithm does not know how many bits will suffice to eliminate, say, a catastrophic cancellation, and assigns infinitely many bits to remain sound.
In this case, Reval falls back to Ziv onion peeling; this is why Reval can require multiple iterations of precision assignment and evaluation before finding a narrow-enough interval. Specifically, Reval uses $\operatorname{slack}[n]$ in place of any infinity in the exponent trick formulas, where $n$ is how many iterations Reval has performed and $\operatorname{slack}[n]$ is a guess as to how many bits might be needed. So, for example, the interval $\bar{z}=[-2^{-93},2^{-108}]$ has $\operatorname{minlog}(\bar{z})=\min(-93,-108)+\operatorname{slack}[n]$. Slack is equal to 512 after the first iteration, and doubles for every further iteration.
This slack mechanism reintroduces a naive precision-doubling algorithm, but only partially:
  • The precisions are still not uniform;
  • The non-zero, non-infinite endpoints still provide useful information about the necessary precision, so even though slack terms involve guesses they are at least somewhat informed; and
  • Slack typically affects only one or a few operations, and the other operations still set their precisions using sound mechanisms.
Because slack only affects a few operations, it can be set more aggressively. The initial guess of 512, for example, is much larger than Sollya’s initial guess of 64.
6.5. Correct rounding
Reval also uses the slack mechanism to achieve correct rounding. Correct rounding requires firmly placing the value to one side or another of a rounding boundary in the target precision. The challenge is that the output can be very close to such a rounding boundary, so that proving it to be on one side or another can require $\bar{z}$ to be much narrower than merely $2^{-p}$. Note that this is rare: in most cases, the two endpoints are either more than 1 ULP apart or match exactly; however, correct rounding is valuable for a variety of use cases (Brisebarre et al., 2024) due to its determinism.
Reval uses the slack mechanism to handle this case: if the output interval rounds to two successive floating-point numbers in the target precision, meaning that Reval has achieved a faithfully- but not correctly-rounded result, the target precision $t$ is increased by $\operatorname{slack}[n]$ before performing precision assignment.
6.6. Other pragmatic tweaks
Rival, the interval library Reval uses, does not seem to provide operators that are tight in the sense of IEEE 1788 (iee, 2015). This means that Equation 3 can produces slightly-too-low precision assignments. In practice, this is not much of a problem, because the interval operators that aren’t tight, like $\sin$ and $\cos$, also use over-approximating exponent trick formulas. But as extra security to avoid issues, Reval uses “${}+5$” instead of “${}+2$” in its precision assignment formulas. This allows for up to 3 ULPs of looseness in each operator, which we believe is more than Rival needs. [^14] We’ve validated this with some experiments, but to our knowledge Rival has not been proven correct and verified bounds are not available. Since much of runtime is spent at very high precisions, the addition of a few bits doesn’t impact performance much.
Reval also supports a number of operators that have no exponent trick formula in Table 2, such as $\operatorname{erfc}$. For these operators Reval simply uses $\operatorname{slack}$ for their $\operatorname{ampl}$ factors. Naturally, Reval’s performance is then no better than naive precision-doubling, but these operators are rare so basic support seems acceptable. For some operators, the slack mechanism is only resorted to for some inputs. For example, $\lceil\log_{2}\operatorname{ampl}_{1}(\operatorname{asin}(\hat{x}_{i}))\rceil\leq\operatorname{maxlog}(\bar{x}_{i})-\operatorname{minlog}(\bar{z}_{i})+1$ as long as $\operatorname{maxlog}(\bar{z}_{i})\leq 0$, but when that condition is not true, slack still has to be added. Similarly, for the rounding operators like $z=\lfloor x\rfloor$, $\lceil\log_{2}\operatorname{ampl}_{z}(\hat{z})\rceil$ is set to $\operatorname{maxlog}(\bar{x})$ if $\hat{z}$ contains more than one integer, $\infty$- ∞ if it contains zero integers, and $\operatorname{slack}[n]$ if it contains exactly one integer (meaning the true value of $x$ might be right near a rounding boundary).
Finally, the $\operatorname{logspan}$ terms in the exponent trick formulas caused issues in early iterations when intervals are very wide. We therefore set $\operatorname{logspan}(\bar{z})=0$; surprisingly, this does not affect soundness, because in the final iteration, where the slack mechanism isn’t used, all values must be known to at least 2 bits of precision, meaning $\operatorname{logspan}(\bar{z})\leq\frac{1}{4}$. Since no exponent trick formula contains more than three $\operatorname{logspan}$ terms, the total added precision from $\operatorname{logspan}$ terms is less then 1. Reval’s use of “${}+5$” instead of “${}+2$” covers this extra bit as well.
7. Evaluation
We investigate the impact of Reval’s precision tuning algorithm through three research questions:
  1. RQ1
    Can Reval evaluate mathematical expressions faster than a uniform-precision baseline?
  2. RQ2
    Can Reval evaluate more mathematical expressions than a uniform-precision baseline?
  3. RQ3
    Does Reval use lower precisions, on average, than a uniform-precision baseline?
We evaluate all three research questions on a set of 528 benchmarks drawn from the Herbie 2.0 benchmark suite, each evaluated at 8256 inputs. For the uniform-precision baseline, we use both Reval with precision tuning disabled and also the state-of-the-art Sollya tool (except for RQ3).151515 Other tools that perform real evaluation include Wolfram Mathematica (Wolfram, 2020), $\lambda_{s}$ (Sherman et al., 2021), the Android and Windows calculators (Boehm, 2020), Calcium (Johansson, [n. d.]), and possibly others. Prior work (Flatt and Panchekha, 2023) has found Mathematica to be much slower than Reval, and the other libraries are not, to our knowledge, optimized for performance.
7.1. Methodology
The 528 benchmarks are mathematical expressions written in the FPCore 2.0 format (Damouche et al., 2016). Benchmarks are short, with 1–59 operations (average 11.6) over 1–9 variables (average 2.5). The most common operations by far are the arithmetic operations, with multiplication at over 30% of operations and addition, subtraction, and division constituting another 50% of operations. The most common non-arithmetic operations are the exponential, power, sine, cosine, logarithm, and square root operations, each approximately 1.5% of operations. The rarest operations were arctangent, hypotenuse, and fused multiply-add. For each benchmark, we randomly sample 8256 input points (that is, values for each variable) using the Herbie 2.0 tool and use the same input points for Reval and both baselines. Input points are sampled using rejection sampling, with inputs that have a domain error according to the baselines or Reval being rejected.
We discard benchmarks with loops, conditionals, data structures (arrays or tensors), or functions not supported by Sollya (mainly cbrt and atan2), which eliminates 17 of the full 547-strong benchmark suite. In addition, two benchmarks (that we know of) trigger a bug in Sollya related to the subexpressions $z^{z}e^{-z}$ and $e^{-z}\ell^{\exp(z)}$, reported to and acknowledged by the developers. We remove these two benchmarks as well. 528 remain, of which 29 are evaluated to single-precision accuracy, while 499 are evaluated to double-precision accuracy; inputs are also given in the same precision. Sollya doesn’t provide hypot and fma functions so these are expanded into $\sqrt{x^{2}+y^{2}}$ and $x\cdot y+z$ during compilation to Sollya’s input language. 98 benchmarks specify a precondition, some of which are simple ranges but others of which involve complex conditionals. Since Sollya does not support preconditions, we do not time precondition evaluation, but use only input points that satisfy the preconditions. Reval is set to correctly round each expression on each of its inputs; however, Sollya instead only attempts to round faithfully.161616 Sollya’s definition of faithful rounding allows it to round a value to any of three possible floating-point values, presumably in the case of exact floating-point results. This is rare but does happen in our evaluation. Moreover, in some cases, the correctly-rounded result is infinite (for example, if the real result is $10^{400}$). In this case, Reval returns infinity but Sollya’s results vary; we track these input points separately and ignore them in RQ1–3.
Baselines
All tools are run with Sollya’s default minimum and maximum precisions or $10+t$ and $(10+t)\cdot 512$, where $t$ is either 24 or 53. These initial and maximum precisions are Sollya’s defaults. For Sollya the necessary precision indicated by setting the prec variable. Sollya’s main evaluation strategy is to evaluate the whole expression under a uniform precision with interval arithmetic, doubling the precision until a faithful rounding is found. However, Sollya does, in various special cases, apply additional tricks, including (from an examination of the source code) Taylor series, l’Hôpital’s rule, and in some cases smaller precision changes than doubling. To the best of our knowledge, however, it does not perform mixed-precision evaluation, except in the case of constants that are cached at high precision. Reval applies the precision tuning algorithm described in this paper, with failure after 5 iterations. We also compare to “Baseline”, which is Reval but applying a uniform precision that doubles with each iteration until the correct rounding has been obtained or the maximum precision has been reached.
For each tool, benchmarks are compiled from FPCore to the tool’s input syntax. The compilations are all straightforward. For Sollya each expression is first defined as a function, with the compilation time not measured, and then that function is evaluated on each input point in turn. The separate compilation step allows Sollya to perform compile-time optimization or to share caches across inputs. Reval also performs a separate compilation step; however, it does not perform any caching except for exact constants like 1. To check that the compilation is performed correctly, we compare the outputs from all three tools (which is how we discovered the previously-mentioned Sollya bug).
Measurement
Evaluation time for an input point must be measured carefully. Reval and Baseline are called within the same process, so a simple nanosecond timer is used. Sollya, however, is started as an external process. To avoid including parsing time and other overhead, Sollya’s time command is used to measure its running time. All tools are subjected to a 20 millisecond timeout. Sollya in particular suffers from what seem like infinite loops; 20 milliseconds was selected as the shortest timeout that still allows Sollya to successfully evaluate most of even the most difficult inputs. A larger timeout would only bias the results toward Reval. Evaluation is performed on a machine with an Intel i7-8700 at 3.7 GHz with 32 GB of DDR4 memory running: Ubuntu 22.04; Sollya 7.0 with MPFI 1.5.3, MPFR 4.1.0, and GMP 6.2.1; and Reval 2.0 with Racket 8.11.1. The evaluation code itself is single-threaded, as are Reval and Baseline and, to our knowledge, Sollya, and the evaluation is run without load on the machine.
7.2. Results
RQ1
Figure 3 shows our results for RQ1. The figure shows evaluations per second for Reval and Baseline, relative to Sollya. Only the 4,359,168 input points where all three tools successfully evaluate the expression are included.171717 This way the handling of invalid inputs, unsamplable inputs, precision limits, and so on do not influence the results. Input points are aggregated by “difficulty”, which is the number of iterations Reval needs to evaluate the point; higher difficulty typically means higher precision. Note that this difficulty metric (chosen because it is easy to compute) is biased against Reval: a point that takes Reval many iterations may nonetheless be easy for Sollya or Baseline. Very roughly, every difficulty level has about 20×20$\times$20 × fewer points than the previous; there are only 30 difficulty-5 points, so its timing results are somewhat unreliable, the the relative standings of the three tools occur repeatedly.
Reval achieves significantly more evaluations per second, both on average and at any difficulty level. Averaged over all points, Reval is 1.25×1.25$\times$1.25 × faster than Sollya and 1.21×1.21$\times$1.21 × faster than Baseline. In fact, Baseline and Sollya perform at similar speeds across the difficulty spectrum, as expected given their similar precision assignment strategy. Reval has the smallest advantage at difficulty 2, which we believe is due to the slack mechanism sometimes overestimating the necessary precision. Reval’s advantage grows as difficulty increases: for the most difficult points, Reval is 2.99×2.$9$ faster than Sollya and 1.71×1.71$\times$1.71 × faster than Baseline. This is expected: more difficult points require higher precisions meaning that precision assignment provides a greater benefit.
Figure 2. Reval’s, Sollya’s, and Baseline’s speed by difficulty of input point. Sollya is normalized to 1.0; higher is better. All three tools perform similarly on the easiest points, and Baseline and Sollya perform similarly throughout, as expected given their very similar core algorithms. However, Reval has a growing advantage on harder inputs, with a speed-up over Sollya of 1.25×1.25$\times$1.25 × overall and 2.99×2.$9$ on the most difficult inputs.
Figure 3. The number of points evaluated by Reval, Sollya, and Baseline, broken down by whether the correctly-rounded result was zero or a non-zero real number. Higher is better. Reval and Baseline both evaluate far more points than Sollya. Reval and Baseline are largely comparable, showing that Reval’s precision tuning does not much affect whether points are successfully evaluated.
RQ2
Figure 3 shows our results for RQ2. This graph considers only points successfully evaluated by at least one, but not all, of the tools. Reval and Baseline evaluate a similar number, with Reval evaluating more points where the correct answer is 0 (3743 to Baseline’s 3741) but fewer points where the correct answer is non-zero (10 to Baseline’s 54). This is expected from precision tuning: Reval can sample more points by avoiding timeouts, but its overapproximations can also cause it to unnecessarily pass the precision limit. In any case, the differences are minimal, showing that precision tuning’s throughput gains come more or less “for free”. Sollya, meanwhile, evaluates only 1635 zero values and 52 non-zero values. We believe the gap is due to Reval and Baseline’s Rival interval arithmetic library handling overflow and underflow in a more rigorous manner (Flatt and Panchekha, 2023) than Sollya’s MPFI. Reval and Baseline also sample 100,346 more points where the correctly-rounded result is infinite; we believe this to be an intentional choice by the Sollya developers, so out of fairness they are not shown in Figure 3.
RQ3
Figure 4 shows our results for RQ3. This plot shows the distribution of execution time by precision across all operations, for both Reval and Baseline. (We were not able to log such granular runtime information in Sollya.) Reval saves a significantly fraction of runtime by executing operations at lower precision than Baseline—note the higher bar at precision $2^{6}$ in exchange for lower bars at precisions $2^{7}$ through $2^{11}$. Precision assignment is also fast, as shown by the separate column on the right: precision tuning consumes 4.1% of run time but saves 17% of run time. Reval spends a tiny fraction of runtime at high precision; in this sense, its precision assignments can be said to be close to optimal.
Figure 4. A weighted histogram showing the distribution of operation execution time by precision, for both Reval and Baseline. Lower is better. Baseline wastes a substantial fraction of its execution time re-computing the same operations at successively higher precisions, leading to a lot of wasted time at precisions between $2^{7}$ and $2^{11}$. By contrast, Reval spends much less time at these high precisions because it can assign high precision only to the few operations that need it.
Other results
Besides the core research questions discussed here, Reval’s precision assignment algorithm provides several additional advantages. One is robustness: on the approximately 72,450 points that none of the three tools could evaluate, Reval terminates much faster. This is because Reval can exit after assigning very high precisions, which takes only one or two iterations. For these points, Reval terminates 11.09×11.09$\times$11.09 × faster than Sollya and 19.87×19.87$\times$19.87 × faster than Baseline. While these points are not considered in RQ1–RQ3, they play a large role in practical deployments of real evaluation, to the extent that the underlying Rival library uses special flags to detect some kinds of unevaluable inputs (Flatt and Panchekha, 2023). Figure 5 breaks down Reval’s greater advantage once unevaluable points are included: early exit allows Reval to skip many useless and expensive high-precision operations, while the uniform precision Baseline has to execute these operations to reach the maximum precision. This also makes it practical for Reval to use a significantly higher maximum precision in practical deployments, which in turn allows it to evaluate more inputs. Finally, compared to Sollya, Reval supports a broader range of operations (cbrt and atan2 are two examples which show up in the Herbie 2.0 benchmark suite), produces correctly- not faithfully-rounded results, and never returns incorrect results, Sollya produces faithfully-, but not correctly-rounded results in 5,000 cases.
Figure 5. A variant of Figure 4, but over all inputs, including those that neither Reval nor Baseline could evaluate. Reval’s early exit allows it to skip most high-precision evaluations, while Baseline wastes a substantial fraction of its runtime uselessly re-evaluating expressions at ever-higher precisions.
7.3. Case Studies
We now examine individual expressions and input points in more detail to highlight aspects of Reval’s precision tuning algorithm.
Slack mechanism
The 2cos benchmark evaluates $f(x,\varepsilon)=\cos(x)-\cos(x+\varepsilon)$ to 53 mantissa bits. When $x$ is big and $\varepsilon$ is small, for example $x=10^{300}$ and $\varepsilon=10^{-300}$, the terms being subtracted differ by a tiny amount, making the output hard to compute. To evaluate the output correctly, the $x+\varepsilon$ term must be computed to at least 600 digits, or about 2000 bits. Then the $\cos$ operations have to be evaluated to approximately 300 digits, or about 1000 bits, and finally the subtraction needs to be evaluated to only 53 bits to achieve a correctly-rounded result.
Reval achieves this precision assignment in three iterations. In the first, untuned iteration, Reval is able to estimate the sizes of $x$, $x+\varepsilon$, and $\cos(x)$. However, because this first iteration uses low precision, $x+e$ covers more than one full period of $\cos$—in fact, it covers approximately $3\cdot 10^{282}$ of them. The interval computed for $\cos(x+e)$ is thus $[-1,1]$, and the subtraction operation likewise has a wide computed interval crossing zero. As a result, during the first tuning pass, Reval applies the slack mechanism twice, once to the subtraction and once to cosine. This results in a precision assignment of $59073+5+\operatorname{slack}[0]=$ 590 for both $\cos$ operations, the slack being introduced by the subtraction, and $590+5+1000+\operatorname{slack}[0]\approx 2100$ for $x+\varepsilon$, where $590$ is the output precision needed, $1000$ is the amplification factor for cosine, and the slack being introduced by the cosine. The amplification factor of $1000$ is a good demonstration of how slack differs from naive uniform precisions. Even though the slack mechanism is used, the $\cos$ operation still provides some information. Specifically, the input to $\cos$ has to be precise enough to at least identify what period of $\cos$ it is in. The full precision assignment is shown in Figure 6.
The first Reval precision assignment:
$\begin{array}[]{llll}P[-]&&&=73\\ P[\cos(x)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil&=73+5+\operatorname{slack}[0]&=590\\ P[\cos(+)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{2}(-)\rceil&=73+5+\operatorname{slack}[0]&=590\\ P[+]&=P[\cos(+)]+5+\lceil\log_{2}\operatorname{ampl}_{1}(\cos(+))\rceil&=590+5+1000+\operatorname{slack}[0]&=2107\end{array}$
The second Reval precision assignment:
$\begin{array}[]{llll}P[-]&&&=73\\ P[\cos(x)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil&=73+5+593+\operatorname{slack}[1]&=1695\\ P[\cos(+)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{2}(-)\rceil&=73+5+593+\operatorname{slack}[1]&=1695\\ P[+]&=P[\cos(+)]+5+\lceil\log_{2}\operatorname{ampl}_{1}(\cos(+))\rceil&=1695+5+1000&=2700\\ \end{array}$
Figure 6. Precision assignment in the first and second iteration for the slack mechanism case study. Note that, even though the slack mechanism partially re-introduces precision doubling, amplification factors still provide useful information.
After re-evaluating with these new precision assignments, the output of the $\cos$ operation no longer crosses zero, but the output of the subtraction still does. The second precision assignment, then, uses the slack mechanism only for the subtraction, assigning a precision of $169573+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil=78+593+\operatorname{slack}[1]=$ 1695 to both $\cos$ operations and a precision of $1695+5+1000=2700$ to the $x+\varepsilon$ term. A third re-evaluation then results in a correctly-rounded result. Note that, while Reval needed only 3 evaluations, a uniform-precision evaluation would require 5 iterations.
Correct rounding mechanism
The placeG benchmark evaluates $f(x,y,z)=(x+y)\times(z+1)$ to 53 mantissa bits. Since the benchmark uses only multiplications and divisions, the exact real result can be represented in floating-point. Consider positive $x$, $y$, and $z$; then all the $\operatorname{ampl}$ and $\operatorname{intro}$ factors are 1 or less. Reval thus assigns low precisions to all operations: 73 bits to the multiplication and 78 to both additions. That works for many positive inputs, but not all.
Some specific inputs suffer from double rounding, such as $x=1.3002052657264033\times 10^{189}$, $y=3.084776002356433\times 10^{188}$, and $z=1\times 10^{-200}$ in double-precision floating-point. These $x$ and $y$ values are exactly one exponent apart, and both end in a 1 bit. This means that their sum, $x+y$, is exactly 54 bits wide, meaning it lies exactly on a rounding boundary. That means it rounds up if $1+z$ is greater than $1$, and down (toward even) if $1+z$ is equal to $1$. In fact, $z+1$ is very slightly greater than 1, which means the answer should round up, but because $z$ is much smaller than $1$, $1+z$ needs to be computed to many, many bits to conclude that it is in fact greater than $1$.
Specifically, in 78 bits, $z+1$ is computed to be $[1,1.000000000000000000000007]$ in the first iteration and the overall expression is computed to be $[1.6...\times 10^{189},1.6...\times 10^{189}]$, where the two endpoints of that interval are successive floating-point numbers. This triggers the double rounding mechanism, causing Reval to evaluate the expression first to $58573+\operatorname{slack}[0]=$ 585 and then to 73+slack[$109773+\operatorname{slack}[1]=$ 1097 bits. This last iteration finally evaluates $1+z$ exactly, proving that it is strictly greater than $1$. By increasing the overall precision demands, the double rounding mechanism was able to eventually determine which side of the rounding boundary the result was on.
Early stopping
The 2nthrt benchmark evaluates $f(x,n)=(x+1)^{1/n}-x^{1/n}$ to double precision. The benchmark is hard to evaluate for large $x$ and small $n$—say, $x=10^{200}$ and $n=10^{-200}$. These inputs require raising a large number to a large power, causing overflow: the intermediate result of $(x+1)^{1/n}$ is $[\approx 2\times 10^{323,228,496},+\infty]$ after the first iteration. The $x^{1/n}$ term likewise overflows, resulting in an output interval of [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] after subtracting the two.
When adjusting precisions for the second iteration, precision tuning algorithm calculates an amplification factor for the subtraction of
[…]
where each term contributes a slack term. This causes Reval to assign a precision of billions of bits to the left-hand term, higher than the maximum allowed precision 32 256. Reval therefore terminates evaluation early.
8. Related Work
H. Boehm was an early advocate of evaluating real-number expressions to high precision using “constructive reals” (Boehm, 1987). A number of implementations of constructive reals exist, including in Java (Boehm, 2004), Python (Briggs, 2006), Caml (Ménissier-Morain, 2005), and others. Perhaps most widely used is the implemenation (Boehm, 2020) in the Android Calculator application. Early implementations used lazy digit sequences, but later work (Lee and Boehm, 1990; Boehm et al., 1986) showed that Ziv’s strategy (Ziv, 1991) with interval arithmetic is typically faster. Interval arithmetic is an old and widely-used technique. Prominent implementations include Arb (Johansson, 2016), Moore (Mascarenhas, 2016), MPFI (Revol and Rouillier, 2005), Rival (Flatt and Panchekha, 2023), and others. Interval arithmetic is used in the Sollya environment (Chevillard et al., 2010) for numerical programming and possibly in Wolfram Mathematica (Flatt and Panchekha, 2023). Many researchers have proposed various tweaks to computable real implementations, such as movability flags to track overflow (Flatt and Panchekha, 2023). Perhaps the most similar to the current work was proposed by Li and Yong (2007): mixed-precision evaluation using “weights”. However, they seem to expect the weights to be user-controlled, whereas Reval derives the precision tuning automatically.
Interval arithmetic and constructive reals are closely related to research into automatic worst-case error bounds for floating-point programs. One important early system was Salsa (Damouche and Martel, 2017; Martel, 2009; Damouche et al., 2015), which used interval arithmetic to bound both ranges of variables and also errors, a strategy adopted by other work as well (Titolo et al., 2018). Rosa (Darulova and Kuncak, 2014) and Daisy (Izycheva and Darulova, 2017) use a similar analysis except with affine arithmetic, which better handles correlated errors. Initially phrased as an abstract interpretation, later papers (Martel, 2018) rephrased the analysis as a type system, where types track both range and precision. FPTaylor (Solovyev et al., 2018) introduced the idea of error Taylor series for automatic worst case error bound estimation, achieving much lower error. Satire (Das et al., 2020) then proposed scaling error Taylor series using abstraction, algebraic simplification, and automatic differentiation, and SeeSaw (Das et al., 2021) extended these techniques to handle bounded control flow. Like Reval, Satire saw higher-order error as a computational bottleneck; unlike Reval, Satire sacrificed soundness for performance, while Reval instead accepted larger error bounds. ATOMU (Zou et al., 2019) proposed computing first-order error from condition numbers. While this provides an algorithmic speedup, ATOMU did not evaluate it; instead, ATOMU focused on condition numbers as a method for driving input generation. In most worst-case error bound tools, the goal is achieving the lowest possible sound error bound, unlike in this paper, where the focus is primarily on computational efficiency.
Automatic precision tuning is also an active area of research, starting with Precimonious (Rubio-González et al., 2013), which used delta-debugging to assign precisions while evaluating accuracy on representative inputs. Later work (Rubio-González et al., 2016) improved scalability to larger programs. Meanwhile, FPTuner (Chiang et al., 2017) and OpTuner (Briggs and Panchekha, 2022) performed sound precision tuning using error Taylor series to derive error models that could be optimized by integer linear programming. This is slower but leads to lower precisions. Finally, POP (Adjé et al., 2021) poses precision tuning as a (non-integer) linear programming problem, allowing it to be solved much more efficiently, though in some settings a “policy iteration” technique is necessary to achieve an optimal precision tuning. The linear programming formulation is quite similar to Reval’s precision tuning algorithm, though no connection is made between the ”amplification factors” and the underlying mathematical operations, and as a result POP does not extend to various transcendental functions.
9. Conclusion
Evaluating real expressions to high precision requires using high-precision interval arithmetic. But state-of-the-art tools use a uniform precision, doubled with every iteration. This wastefully allocates too much precision to non-accuracy-critical operations, and can also overestimate precision or require too many iterations to find the right precision. We instead introduce a fast, sound method for assigning precisions to interval operators so as to correctly-round a real expression. Our method is based on a variation on error Taylor series and condition numbers, called amplification factors, which is sound, computable quickly, and specialized to interval arithmetic. A key optimization is the use of “exponent tricks”, which allow computing amplification factors using only operations on machine integers. As a result, Reval is able to evaluate real expressions 1.25×1.25$\times$1.25 × faster than the state-of-the-art Sollya tool across our suite of 528 benchmarks, with Reval’s advantage rising to 2.99×2.$9$ for the most difficult input points. It achieves this speedup largely by using a sophisticated mixed-precision assignment algorithm that evaluates operations at high precision only when necessary.

### v2

@@
-Fast Real Evaluation Through Sound Mixed-Precision Tuning
+Fast Sound Error Bounds for Mixed-Precision Real Evaluation
 Abstract.
-Evaluating a real-valued expression to high precision is a key building block in computational mathematics, physics, and numerics.
+Evaluating real-valued expressions to high precision is a key building block in computational mathematics, physics, and numerics.
 A typical implementation uses a uniform precision for each operation, and doubles that precision until the real result can be bounded to some sufficiently narrow interval.
@@
 To do so, we introduce a sound variation of error Taylor series and condition numbers, specialized to interval arithmetic, that can be evaluated with minimal overhead thanks to an “exponent trick”.
-Our implementation, Reval, achieves a speed-up of 1.25×1.25$\times$1.25 × compared to the state-of-the-art Sollya tool, with the speed-up increasing to 2.99×2.$9$ on the most difficult input points.
+Our implementation, Reval, achieves an average speed-up of 1.47×1.47$\times$1.47 × compared to the state-of-the-art Sollya tool, with the speed-up increasing to 4.92×4.92$\times$4.92 × on the most difficult input points.
 An examination of the precisions used with and without precision tuning shows that the speed-up results come from quickly assigning lower precisions for the majority of operations.
-Interval arithmetic, precision tuning, correct rounding, rounding error, error Taylor series, condition numbers, arbitrary precision
+Interval arithmetic, precision tuning, correct rounding, rounding error, error Taylor series, condition numbers, arbitrary precision.
 copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXX
 1. Introduction
-Many tasks in computational science require computing real expressions to high precision, including computational number theory (Hakert et al., 2024; Bailey, 2001), physics (Bailey and Borwein, 2015), geometry (Seel, 2024), and commutative algebra (Grayson and Stillman, [n.
+Many tasks in computational science require computing real expressions to high precision, including computational number theory (Hakert et al., 2024; Bailey, 2001), physics (Bailey and Borwein, 2015), geometry (Seel, 2024; Sherman et al., 2019), and commutative algebra (Grayson and Stillman, [n.
 d.]).
-Moreover, computing real expressions to high precision is also a computational bottleneck in the Remez or Chebyshev approximation algorithms (Driscoll et al., 2014; Remez, 1934; Lim et al., 2021); in numerical compilers like Herbie (Panchekha et al., 2015; Flatt and Panchekha, 2023); in implementations of math library functions like $\exp$ and $\operatorname{pow}$ (Bugzilla, 2006; archive of the discontinued LLVM Phabricator instance, 2020); and even in simple desktop calculators such as the one available on Android (Boehm, 2020).
-Computing real expressions to high precision is challenging because computer arithmetic is subject to rounding error.
+Real expressions evaluation is also necessary for the Remez and Chebyshev approximation algorithms (Driscoll et al., 2014; Remez, 1934; Lim et al., 2021); numerical compilers like Herbie (Panchekha et al., 2015; Flatt and Panchekha, 2023); and math library functions like $\exp$ and $\operatorname{pow}$ (Bugzilla, 2006; archive of the discontinued LLVM Phabricator instance, 2020).
+For example, to synthesize a math library function using RLibm (Lim et al., 2021) one first needs a correctly-rounded evaluation of the function on all 4 billion-odd 32-bit floating-point inputs.
+Naturally, both performance and generality are critical.
+The challenge is that computer arithmetic is subject to rounding error.
 Real number evaluation thus typically uses high-precision interval arithmetic (Chevillard et al., 2010; Panchekha et al., 2015; Johansson, 2016), in particular a strategy known as Ziv’s onion-peeling method (Ziv, 1991).
-This method first evaluates the real expression using interval arithmetic and checks if the resulting interval is narrow enough for whatever purpose the user may have in mind.
-If it’s not, Ziv’s strategy repeatedly re-evaluate the same expressions at higher and higher precision until a narrow-enough output interval is achieved.
-A typical implementation uses a single, uniform precision for evaluating the whole expression and doubles that precision with every re-evaluation.
-However, evaluating the whole expression at a uniform precision is wasteful and slow: it forces us to evaluate the whole expression at the precision needs of its highest-precision operation.
-A typical example might be evaluating $\sin(\pi\cdot x)+2$ for a very large number $x\approx 10^{100}$.
-To compute an answer to, say, 20 decimal places, the multiplication by $\pi$ must be performed at extremely high precision, producing at least 120 correct output digits.
-However, the rest of the computation doesn’t need nearly this many bits: the output of $\sin$ is 0 and doesn’t need any digits, and the addition needs to produce only 20 correct digits.
-Using a uniform precision for this example would waste time and memory.
-And when a high precision is needed, the uniform precision-doubling approach must perform many wasteful iterations in order to discover the necessary high precision, which it might overestimate by a factor of two.
-A better approach would tune the precision of each operation individually.
-We implement just that in Reval, a new library for real evaluation using intervals.
+This method evaluates the real expression using interval arithmetic and checks if the resulting interval is narrow enough for whatever purpose the user may have in mind.
+If it’s not, Ziv’s strategy re-evaluates at higher and higher precision until it is.
+A key parameter in Ziv’s strategy is the precision to evaluate the expression at.
+The standard implementation uses a single, uniform precision for all operations that doubles with every re-evaluation.
+This choice is widely understood to be sub-par.
+The implementation of the Marshall language (Sherman et al., 2019), for example, suggests that “we should do something more intelligent about that (not all subexpressions of $e$ need the same precision)”, while the Sollya language (Chevillard et al., 2010) implementation describes the choice of precision as a “matter of religion” with doubling being “fine in practice.” Using a uniform precision for all operations means using a precision that is too high for all but the most precision-sensitive operations, wasting valuable time.
+Nonetheless, no better, practical method currently exists.
+We provide one in Reval, a new library for real evaluation using intervals.
 [^1] The name is also a pun on Rival, the underlying interval library Reval uses.
 Reval uses a precision assignment algorithm to determine what precision to use for every operation in the program.
-The computed precision assignment is based on sound worst-case error bounds and can be computed quickly, with minimal overhead, and results in dramatically lower precision needs for the overwhelming majority of operations in an expression.
-This allows Reval to evaluate real expressions to high precision faster than state-of-the-art libraries such as Sollya (Chevillard et al., 2010).
-Reval derives a sound floating-point error bound using an algorithm inspired by error Taylor series and condition numbers as used in FPTaylor (Solovyev et al., 2018), Satire (Das et al., 2020), and ATOMU (Zou et al., 2019).
-However, Reval’s error bounds are specialized to interval arithmetic, which allows for sound bounds without the need for a “higher-order error” term that requires second derivatives.
-Because Reval’s bounds only require first derivatives, they can be computed in time linear to the size of the expression, instead of the quadratic time required for “higher-order error” terms.
-To avoid slow arbitrary-precision computations, Reval uses an “exponent trick” which allows it to over-approximate the error bound using only operations over machine integers.
-The resulting error bounds can then be used for precision assignment, yielding sound, reasonably tight error bounds computed with minimal overhead.
-We evaluate Reval on 4,359,168 computations (528 benchmarks each evaluated on 8 256 inputs) drawn from the Herbie 2.0 benchmark suite, and compare it to the state-of-the-art Sollya tool (Chevillard et al., 2010).
-Reval is 1.25×1.25$\times$1.25 × faster while successfully evaluating a similar number of computations, with an even larger speedup of 2.99×2.$9$ on the most difficult computations.
-A comparison with uniform precision-doubling shows that Reval’s performance advantage is due to executing most operations at substantially lower precision.
-Besides these quantitative measures, Reval’s precision assignment algorithm is also more robust to invalid inputs.
+Critically, each operation gets its own precision.
+This leads to dramatic speed-ups because most operations are executed at low precision even if a few require very high precisions.
+Reval’s approach also has secondary benefits.
+The standard precision-doubling approach can be thought of as an exponential search for a high-enough precision; because Reval computes the correct precision analytically, it can typically skip the search process and jump straight to the correct precision.
+It can also avoid recomputing operations where a good-enough answer is already known, avoid over-estimating due to discretization, and discard erroneous inputs more quickly.
+This allows Reval to evaluate real expressions faster than state-of-the-art libraries such as Sollya (Chevillard et al., 2010).
+Reval’s precision assignments are based on sound worst-case rounding error bounds and can be computed quickly, with minimal overhead.
+Following prior work (Solovyev et al., 2018; Das et al., 2020; Zou et al., 2019), these bounds are based on error Taylor series and condition numbers.
+However, Reval recognizes that, for interval arithmetic, sound techniques can be dramatically faster, because interval arithmetic does require computing an expensive “higher-order error” term.
+This allows Reval to compute error bounds, and thus precision assignments, in time linear to the size of the expression, instead of the quadratic (or worse) time of prior sound tools.
+Moreover, our implementation also uses an “exponent trick” to avoid arbitrary-precision operations during precision tuning, leading to further speed-ups.
+The resulting error bounds are sound, reasonably tight, and computed with minimal overhead.
+We evaluate Reval on 567,238 computations drawn from 493 benchmarks in the Herbie 2.0 benchmark suite, and compare it to a uniform-precision baseline and the state-of-the-art Sollya tool (Chevillard et al., 2010).
+Reval is 1.62×1.62$\times$1.62 × faster, with speedups of up to 4.92×4.92$\times$4.92 × on the most difficult computations.
+We show that this performance advantage is due to the highly non-uniform precisions assigned by Reval.
+Reval’s approach also has additional benefits, such as faster convergence, fewer operator evaluations, and better handling of invalid inputs, all of which contribute to Reval’s performance advantage.
 In summary, this paper contributes:
-  • A sound error bound algorithm specialized to interval arithmetic (Section 4).
+  • A fast, sound error bound algorithm specialized to interval arithmetic (Section 4).
   • An “exponent trick” for computing these sound error bounds efficiently (Section 5).
@@
 2. Overview
-[…]
+Consider the task of evaluating the expression $(1.0-\cos(x))/\sin(x)$ for small values of $x$, such as [^2] Reval inputs are MPFR values, so $x$ is actually equal to the double-precision number closest to $10^{-12}$.
+$10^{-12}$.
+Floating-point arithmetic suffers from accuracy problems: in double-precision floating-point, this expression evaluates to $0.0$, which must be incorrect because $\cos(x)<1$.
+Since double precision isn’t enough, a natural option is arbitrary-precision floating-point, which can be accessed via libraries like MPFR.
+But this raises its own question: what precision?
+With 53 bits of precision—MPFR’s analog of double precision—this expression returns $0.0$, and it returns $0.0$ at 70 bits of precision as well.
+But at 80 bits of precision, it returns $8.27\dotsb 10^{-13}$, at 85 bits $4.91\dotsb 10^{-13}$, and at 90 bits $5.00\dotsb 10^{-13}$.
+None of these are the right answer, or even particularly close: at 90 bits the relative error is $4\cdot 10^{-5}$.
+Intuitively, a high-enough precision should work—with 5000 bits of precision, the relative error is vanishingly small—but it’s hard to know how much precision is enough, and high precisions are much more expensive to compute with.
 For this reason, real evaluation is typically done using interval arithmetic, which can certify that a computed answer is actually correct.
@@
 When subtracted from $1.0$, the resulting interval is approximately $[4.13\cdot 10^{-25},5.17\cdot 10^{-25}]$, and the final interval is $[4.13\cdot 10^{-13},5.17\cdot 10^{-13}]$.
-Note that this final interval includes the correct result—but is too wide to actually identify it.
-A higher precision, such as 160 bits, will compute a narrower interval like
+This interval includes the correct result but is too wide to actually identify it.
+A higher precision, such as 160 bits, will compute a narrower final interval like
 […]
@@
 In Reval (right), precision tuning is performed if a correct rounding is not found, deriving a different precision for each operation.
-This overall strategy—evaluating in arbitrary-precision interval arithmetic, and raising the precision until the interval is narrow enough—is the standard approach to accurately evaluating real expressions, diagrammed in Figure 1.
+This overall strategy—evaluating in arbitrary-precision interval arithmetic, and raising the precision until the interval is narrow enough—is the standard “Ziv’s method” for accurately evaluating real expressions, diagrammed in Figure 1.
 The key challenge is choosing the precision.
-If the goal is 53 correct bits—a correct double-precision result—trying a uniform 80- and 160-bit evaluation is wasteful.
-For example, 160 bits are only actually necessary for $\cos(x)$; all other operations could be evaluated in as few as 57 bits and still get the right answer.
-Moreover, doubling precision on each iteration can overshoot: if we’d started at 64 instead of 80 bits, then we’d try 128, find too-wide an interval, and then have to evaluate every operation at 256 bits.
-All of this inefficiency comes back to the fact that we are, fundamentally, guessing what precision will be sufficient.
-It would be better to derive the precision from some analytical method.
+If the goal is 53 correct bits—a correct double-precision result—a uniform 160-bit evaluation is wasteful; only $\cos(x)$ needs to be evaluated at 160 bits, while all other operations could be evaluated at 57 bits and still get a narrow-enough final interval.
+More generally, the best precision assignment is typically highly non-uniform; the question is just how to find it.
 That is exactly what Reval does.
-Suppose we wish to compute $(1-\cos(x))/\sin(x)$ at $x=10^{-12}$, as above, to double precision.
-Reval first evaluates every operation at a fixed low precision.
-It then uses the computed intervals to compute amplification factors for each operation.
-Intuitively, the amplification factor describes how much the operation amplifies error in its inputs.
-For example, the subtraction operation has an amplification factor of $2^{81}$ because it cancels the first 81 bits of its two inputs.
-These amplification factors are related to condition numbers and reverse-mode automatic differentiation, though with slight variations to make them sound.
-The amplification factor (specifically, its logarithm) then indicates how many more bits the inputs of an operator must have for the output to have some target precision.
-Amplification factors are computed using “exponent tricks”; the subtraction operation’s amplification factor is computed by simply subtracting the unbiased exponents of the high endpoint of $\cos(x)$ and the low endpoint of $1-\cos(x)$.
-[^3] Plus 1 to account for the significands.
-Subtracting machine integers, and Section 5 provides similar exponent tricks for other common operations (Section 5).
+After an initial evaluation at low precision, Reval computes amplification factors for each operation.
+These rigorously bound how each operation amplifies error in its inputs.
+For example, the subtraction operation has an amplification factor of $2^{81}$ because it cancels the first 81 bits of its two inputs, while the division operation has an amplification factor of $1$.
+These amplification factors are related to condition numbers and reverse-mode automatic differentiation and are computed using “exponent tricks”; basically, they are based on rigorous and sound worst-case rounding error bounds.
+The theory is laid out in Section 4.
 The amplification factor is then used to assign a precision to each subexpression.
@@
 of $16483+\log_{2}1+\log_{2}2^{81}=$ 164.
-This precision assignment is not a guess: it is guaranteed to yield a narrow-enough output interval.
-So, after one more evaluation with this new precision assignment, Reval computes the correctly-rounded double-precision result.
-Deriving a sufficient precision assignment is not always so easy.
+Because amplification factors come from worst-case error bounds, this precision assignment is not a guess: it is guaranteed to yield a narrow-enough output interval.
+So, after re-evaluation with this new precision assignment, Reval computes the correctly-rounded double-precision result.
+In rare cases, amplification factors cannot be computed.
 For example, consider the input $x=10^{-100}$ to the same expression.
 In its first evaluation at the same fixed precision, Reval computes an interval for the subtraction of $[0,1.033\cdot 10^{25}]$.
-Because this interval includes 0, Reval cannot bound the exponent of this value, which in turn makes the exponent trick formula return a useless bound of infinity.
+Because this interval includes 0, Reval cannot bound the exponent of this value, which in turn produces a useless amplification factor of infinity.
 Reval works around this by guessing $84+512=596$ for the amplification factor (where $512$ is a fixed constant) and ends up assigning 679 bits of precision to the $\cos(x)$ operation.
 All other operations, however, need no guesses and end up with the starting precision.
-Reval now performs a second evaluation at these new precisions.
-This second evaluation does not yield a narrow-enough interval: Reval’s guess was insufficient.
-Reval thus re-performs precision assignment.
-This time the interval for the subtraction is $[4.999\cdot 10^{-101},5.001\cdot 10^{-101}]$ and Reval can compute the amplification factor to be $2^{666}$.
-Reval then assigns a precision of 749 bits to the $\cos(x)$ term, and re-evaluating again finally yields a correctly-rounded result.
+Re-evaluation with these precisions is now not guaranteed to yield a narrow-enough interval, due to the guess, but it does produce a narrow interval for the subtraction, $[4.999\cdot 10^{-101},5.001\cdot 10^{-101}]$, which allows Reval to compute an amplification factor in a second precision-tuning pass.
+The second re-evaluation then yields a correctly-rounded result.
 To summarize, Reval initially performs an evaluation at some fixed low precision.
@@
 In most cases, the precision assignment guarantees that, after re-evaluation, a narrow-enough interval will be found.
-[^5] Narrow enough for a “faithful”, though possibly not “correct”, rounding (Brisebarre et al., 2024); see Section 6 However, in other cases, Reval cannot bound the amplification factors and is forced to guess; in this case, it makes guesses with successively larger precisions until a sufficient precision assignment can be computed.
+[^4] Narrow enough for a “faithful”, though possibly not “correct”, rounding (Brisebarre et al., 2024); see Section 6 However, in other cases, Reval cannot bound the amplification factors and is forced to guess; in this case, those guesses can at least allow it to compute amplification factors in later iterations.
 The rest of this paper is organized as follows: Section 3 covers in brief floating-point and interval arithmetic, rounding, and Reval’s mathematical foundations, error Taylor series and condition numbers.
@@
 A normal floating-point number is a number ±$\pm(1+m)2^{e}$ where the mantissa $m$ is a value in $[0,1)$ represented with some fixed precision $p$ and where the exponent $e$ is a signed integer.
-Floating-point numbers also include two zeros (positive and negative), two infinities, a Not-a-Number or NaN value, and a separate class of subnormal values, but in this paper we are largely concerned with normal numbers as other techniques (Flatt and Panchekha, 2023) are available for handling under-flow, over-flow, and the corresponding zero and infinite values.
+There are also various special cases, but other techniques (Flatt and Panchekha, 2023) are available for handling them, and they are not a focus for this paper.
 We are specifically concerned with arbitrary-precision floating-point arithmetic as implemented in the well-known MPFR library (Fousse et al., 2007).
@@
 A real number $x$ can be rounded to a floating-point value (at precision $p$).
-If the real number is already exactly representable, that is easy, but typically $x$ lies strictly between two floating-point values, in which case we say that either one is a faithful rounding of $x$.
-A faithful rounding of $x$, $\hat{x}$, is typically different from $x$, but at precision $p$ the relative error $(x-\hat{x})/x$ between the two is at most $2^{-p}$, assuming $x$ is within the normal exponent range.
+If the real number is already exactly representable, that is easy, but typically $x$ lies strictly between two floating-point values, in which case we say that either one is a faithful rounding $\hat{x}$ of $x$.
 A specific rounding mode, such as the default round-to-nearest ties-to-even, may prescribe one of those two values as the correct rounding of $x$.
+Correct rounding is the stricter notion, since most real numbers have two faithful roundings but only one correct rounding; for this reason correct rounding is required in many applications (Brisebarre et al., 2024).
 A floating-point implementation $\hat{f}$ of some mathematical function $f$ is said to be faithfully rounded if $\hat{f}(\hat{x})$ is a faithful rounding of $f(\hat{x})$ for all floating-point inputs $\hat{x}$ in its domain, and likewise for correct rounding.
-Correct rounding is the stricter notion, since most real numbers have two faithful roundings but only one correct rounding.
-Moreover, correct rounding is deterministic, while faithful rounding is not.
-For this reason, Reval aims for correctly-rounded results.
 A significant advantage of faithful rounding is its relationship with floating-point error.
-If a number $x$ can be bounded to some range $[x_{1},x_{2}]$, and if the relative error of $x_{1}$ and $x_{2}$ is less than $2^{-p}$, then a faithful rounding of $x$ can be determined.
+If a real number $x$ can be bounded to some range $[x_{1},x_{2}]$ with a relative error of $x_{1}$ and $x_{2}$ less than $2^{-p}$, then a faithful rounding of $x$ can be determined.
 The same is not true of correct rounding: even a very narrow interval $[x_{1},x_{2}]$, much narrower than $2^{-p}$, might still contain the rounding boundary where the correct rounding switches from rounding up to rounding down.
@@
 Such an interval is called valid (iee, 2015); the narrowest such interval is called tight.
-Interval arithmetic is just an abstract domain over floating-point numbers, and the condition $y_{\operatorname{lo}}\leq f(x)\leq y_{\operatorname{hi}}$ is just soundness of the transfer function $\bar{f}$.
-Many implementations of interval arithmetic exist, both over hardware floating-point numbers (such as Gaol (Goualard, 2019)) and over arbitrary-precision floating-point (such as Rival (Flatt and Panchekha, 2023), used by Reval, or MPFI (Revol and Rouillier, 2005), used by Sollya).
-These implementations guarantee validity but may depart from tightness in certain difficult cases.
+Interval arithmetic is an abstract domain over floating-point numbers, and validity is just soundness of the transfer function $\bar{f}$.
+Many implementations of interval arithmetic exist, both over hardware floating-point numbers (Goualard, 2019) and over arbitrary-precision floating-point (Flatt and Panchekha, 2023; Revol and Rouillier, 2005).
+These implementations guarantee validity but depart from tightness for certain difficult inputs.
 The benefit of interval arithmetic is that intervals are computed using floating-point arithmetic but bound the behavior of real numbers.
@@
 The endpoints $y_{p,\operatorname{lo}}$ and $y_{p,\operatorname{hi}}$ are then rounded outward to precision $t$.
-If the two results are successive floating-point numbers in precision $t$, then either one is a faithful rounding of the expression.
-Otherwise, the same process can be repeated at higher and higher working precisions $p$.
+If the two results are successive floating-point numbers in precision $t$, then either one is a faithful rounding of the expression; if the two results are equal, then they are a correct rounding.
+Otherwise, this can be repeated at higher and higher working precisions $p$ to obtain a faithful or correct rounding, as required.
 Error Taylor Series
@@
 If the $\hat{f}_{i}$ are faithfully or correctly rounded then we can write $\hat{z}_{i}=f_{i}(\hat{x}_{i},\hat{y}_{i})(1+\varepsilon_{i})$ for some $|\varepsilon_{i}|<2^{-p}$.
-[^7] More care is needed if the computation of $e$ can be subject to under- or overflow errors.
-However, this paper does not consider these errors, as prior work (Flatt and Panchekha, 2023) has explained how to handle them in the context of evaluating real expressions.
-If we perform this rewrite for every operation in the expression, then the value of $\hat{z}$ becomes a real-valued function $\hat{z}(\vec{\varepsilon})$ of many $\varepsilon_{i}$.
+[^6] More care is needed to handle under- or overflow errors, which in Reval are handled by the underlying interval library (Flatt and Panchekha, 2023).
+If we apply this bound to every $\hat{z}_{i}$, $\hat{z}$ becomes a real-valued function $\hat{z}(\vec{\varepsilon})$ of many $\varepsilon_{i}$.
 A real evaluation of $\hat{z}$ then corresponds to $\hat{z}(\vec{0})$, since in a real evaluation there is no rounding error.
@@
 This expansion is
-$\hat{z}(\vec{\varepsilon})=\hat{z}(\vec{0})+\sum_{\varepsilon}\varepsilon\frac{\partial\hat{z}}{\partial\varepsilon}(\vec{0})+O(\varepsilon^{2})$
-The first term is the real evaluation; the second term is called the “first-order error”; and the third term is called the “higher-order error”.
-For example, for the overview example $(1-\cos(x))/\sin(x)$, the first-order error is
-$-\frac{\cos(x)}{\sin(x)}\varepsilon_{\cos}+\frac{1-\cos(x)}{\sin(x)}\varepsilon_{-}-\frac{1-\cos(x)}{\sin(x)}\varepsilon_{\sin}+\frac{1-\cos(x)}{\sin(x)}\varepsilon_{/}$
-More generally, both the first- and second-order errors can be computed by taking symbolic derivatives of $\hat{z}$.
-They can then be bounded using a global non-linear optimizer for first-order error and interval arithmetic for second-order error.
+$\hat{z}(\vec{\varepsilon})=\hat{z}(\vec{0})+\sum_{\varepsilon}\varepsilon\frac{\partial\hat{z}}{\partial\varepsilon}(\vec{0})+O(\varepsilon^{2}).$
+The first term is the real result; the second term is called the “first-order error”; and the third term is called the “higher-order error”.
+The first- and second-order errors can be computed by taking symbolic derivatives of $\hat{z}$ and then bounding them using a global non-linear optimizer for first-order error and interval arithmetic for second-order error.
 The two terms use different optimization techniques because the higher-order error is both much smaller—order $\varepsilon^{2}$ instead of $\varepsilon$—and also has quadratically more terms, meaning a faster, looser bound is preferred.
@@
 Condition numbers simplify the computation of first-order error; while we first learned of the connection in ATOMU (Zou et al., 2019), the basic insight is the same as that in reverse-mode automatic differentiation.
-Suppose one wishes to bound the relative error of a floating-point vs real computation of an expression $\hat{z}$.
-Then, following the error Taylor series idea above, one needs to compute the first-order error divided by $\hat{z}$, which involves computing
-$\frac{(\partial\hat{z}/\partial\varepsilon_{i})(\vec{0})}{\hat{z}(\vec{0})}.$
+Suppose one wishes to bound the relative error, not absolute error, using error Taylor series.
+One then needs to compute the first-order error divided by $\hat{z}$, which involves computing $(\partial\hat{z}/\partial\varepsilon_{i})(\vec{0})/\hat{z}(\vec{0})$.
 Condition numbers rearrange this computation into a sum of products
-$\sum_{p}\prod_{f_{j}\in p}\Gamma_{f_{j}}.$
-Here, the sum ranges over all paths $p$, where a path means a path in the expression tree from the root to some intermediate value.
+$\sum_{p}\prod_{f_{j}\in p}\Gamma_{f_{j}},$
+where the sum ranges over all paths $p$ from the root of the expression tree to some intermediate node.
 For example, in the overview example $(1-\cos(x))/\sin(x)$, the path from the root to $\cos(x)$ involves a division and a subtraction, while the path from the root to $\sin(x)$, involves just a division.
-Then the $f_{j}$ ranges over each operation in that path, and for each path their condition numbers $\Gamma_{f_{j}}$ are multiplied together.
+Then $f_{j}$ ranges over each operation in that path, and the condition numbers $\Gamma_{f_{j}}$ of those operations are multiplied together and summed across paths.
 The condition number $\Gamma_{f}$ is equal to $xf^{\prime}(x)/f(x)$, which can be evaluated symbolically for all operations $f_{i}$ in the expression $\hat{z}$.
@@
 Reval’s precision assignment algorithm is based on error bounds for floating-point computations.
-At a high level, these error bounds are based on error Taylor series and condition numbers, with some variations to ensure soundness and performance.
-Their derivation is fairly technical; we invite readers uninterested in the technical details to focus on the high-level description in Section 4.1 and then on Equations (1), (2), and (3), which are the key results proved in the rest of this section.
-4.1. High-level Approach
 Intuitively, the endpoints computed in interval arithmetic bound all possible faithfully-rounded evaluations of a given expression.
@@
 If that error bound is parameterized by the precision of each operation, then the sound error bound is an inequality that we can solve for an appropriate precision assignment.
-Importantly, these error bounds need to be sound, always over-approximating the error; otherwise, a too-low precision will be assigned and the resulting output interval will be too wide.
-But if the error is over-approximated by too much, too-high a precision will be assigned, which will make evaluation take too long.
+Ideally the bounds are not only sound but also relatively tight, to avoid over-estimating the necessary precision.
 The need for soundness and accuracy suggests error Taylor series, but performance is a problem.
-Error Taylor series are only sound if the “higher-order error” term is included, and that term is expensive to compute: it requires computing quadratically-many second derivatives of the expression.
-We instead introduce a novel variation on error Taylor series that achieves soundness using only first derivatives.
-Equation 1 shows that variation.
-Even this variation, computed naively, would take $O(n^{2})$ time to compute (where $n$ is the number of real operations), but an algebraic reorganization inspired by (and related to) condition numbers allows computing it in $O(n)$ time.
-That reorganization is shown in Equation 2.
-Finally, the error bound can be solved for an appropriate precision assignment; the solution is shown in Equation 3.
+Error Taylor series are only sound if the “higher-order error” term is included, which involves the quadratically-many second derivatives of the expression.
+We introduce a slight variation on error Taylor series that achieves soundness using only first derivatives; Equation 1 shows that variation.
+An algebraic reorganization inspired by (and related to) condition numbers allows computing this variation in $O(n)$ time; Equation 2 shows that reorganization.
+Finally, the error bound can be solved for an appropriate precision assignment; Equation 3 shows that solution.
+The derivations of each equation are fairly technical and we invite readers to skip them and jump directly equations on their first read-through.
 4.1. Sound First-Order Error
@@
 Each operation also has a corresponding real-number operation $z_{i}=f(x_{i},y_{i})$.
-We also assume each $\hat{f}_{i}$ is faithfully rounded,888 Reval’s interval arithmetic is not tight, especially in challenging cases like $\cos(\pi)$, but adding a couple of extra bits allows for a couple of ULPs of looseness; see Section 6 for more.
+We also assume each $\hat{f}_{i}$ is faithfully rounded,777 Reval’s interval arithmetic is not tight, especially in challenging cases like $\cos(\pi)$, but adding a couple of extra bits allows for a couple of ULPs of looseness; see Section 6 for more.
 so that
 $\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})=f_{i}(\hat{x}_{i},\hat{y}_{i})(1+\varepsilon_{i}),$
-for some $|\varepsilon_{i}|\leq 2^{-p_{i}}$, where $p_{i}$ is the precision used for this operation.999 This formula is not correct in the presence of under- and over-flow, though many cases of such are handled using movability flags (Flatt and Panchekha, 2023), a feature of the interval library we use.
+for some $|\varepsilon_{i}|\leq 2^{-p_{i}}$, where $p_{i}$ is the precision used for this operation.888 As before, under- and over-flow are handled using movability flags (Flatt and Panchekha, 2023), a feature of the interval library we use.
 The remaining cases are handled by the slack mechanism of Section 6.
@@
 Notice that $\hat{z}_{i}(\vec{0})=z_{i}$; the numerator of the relative error formula is thus $\hat{z}_{i}(\vec{\varepsilon})-\hat{z}_{i}(\vec{0})$.
-But $\hat{z}_{i}(\vec{0})$ is the first term of the Taylor series for $\hat{z}_{i}$.
-Here we diverge from error Taylor series, applying the Lagrange remainder theorem at order 1 instead of order 2: [^10] There are many variations of the Lagrange remainder theorem; the statement below is one of them but can also be seen as a corollary of the mean value theorem.
+But $\hat{z}_{i}(\vec{0})$ is the first term of the Taylor series for $\hat{z}_{i}$, so we are looking to bound the error of a Taylor series approximation.
+To do so, we apply the Lagrange remainder theorem at order 1: [^9] There are many variations of the Lagrange remainder theorem; the statement below is one of them but can also be seen as a corollary of the mean value theorem.
 $\hat{z}_{i}(\vec{\varepsilon})-\hat{z}_{i}(\vec{0})=\sum_{j\leq i}\varepsilon_{j}\frac{\partial\hat{z}_{i}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})$
 for some $0\leq\vec{\varepsilon^{*}}\leq\vec{\varepsilon}\leq 2^{-p_{i}}$ (or the symmetrical bound for $\varepsilon_{j}<0$).
-In this bound, $\vec{\varepsilon}$ represents the rounding error of a faithfully-rounded execution, so $\vec{\varepsilon^{*}}$ represents the rounding error of some other, also faithfully-rounded execution (of $\partial f/\partial\varepsilon_{j}$).
-We don’t know the signs of the $\varepsilon_{i}$, just a bound on their magnitude, but using the triangle inequality we can still bound the relative error of $\hat{z}_{i}$:
+In this bound, $\vec{\varepsilon}$ represents the rounding error of a faithfully-rounded execution, so $\vec{\varepsilon^{*}}$ represents the rounding error of some other, also faithfully-rounded execution.
+Since we don’t know the signs of $\hat{\varepsilon}$, we bound the relative error of $\hat{z}_{i}$ using the triangle inequality:
 (1) $\left|\frac{\hat{z}_{i}-z_{i}}{z_{i}}\right|=\left|\frac{\hat{z}_{i}(\vec{\varepsilon})-\hat{z}_{i}(\vec{0})}{z_{i}}\right|\leq\max_{\vec{\varepsilon^{*}}}\sum_{j\leq i}2^{-{p_{j}}}\left|\frac{1}{z_{i}}\frac{\partial\hat{z}_{i}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|.$
@@
 $+(1+2^{-p_{i}})\left(\max_{\vec{\varepsilon^{*}}}\left|\frac{x_{i}(\partial_{1}f_{i})(x^{*}_{i},y^{*}_{i})}{f_{i}(x_{i},y_{i})}\right|\right)\left(\max_{\vec{\varepsilon^{*}}}\sum_{j<i}2^{-p_{j}}\left|\frac{1}{x_{i}}\frac{\partial\hat{x_{i}}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|\right)$
-$+(1+2^{-p_{i}})\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\left|\frac{y_{i}(\partial_{2}f_{i})(x^{*}_{i},y^{*}_{i})}{f_{i}(x_{i},y_{i})}\right|\right)}_{\operatorname{ampl}_{2}(\hat{z}_{i})}\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\sum_{j<i}2^{-p_{j}}\left|\frac{1}{y_{i}}\frac{\partial\hat{y_{i}}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|\right)}_{\operatorname{bound}(\hat{y}_{i})}$
+$+(1+2^{-p_{i}})\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\left|\frac{y_{i}(\partial_{2}f_{i})(x^{*}_{i},y^{*}_{i})}{f_{i}(x_{i},y_{i})}\right|\right)}_{\operatorname{ampl}_{k}(\hat{z}_{i})}\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\sum_{j<i}2^{-p_{j}}\left|\frac{1}{y_{i}}\frac{\partial\hat{y_{i}}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|\right)}_{\operatorname{bound}(\hat{y}_{i})}$
 where $x^{*}_{i}$ and $y^{*}_{i}$ are $\hat{x}_{i}$ and $\hat{y}_{i}$ evaluated at $\vec{\varepsilon^{*}}$. Rewriting this gargantuan formula using the helper functions defined by the braces, we get:
@@
 However, condition numbers bound only the first-order error so are unsound, while Equation 2 is sound.
-This in turn causes differences like the $\operatorname{intro}(\hat{z}_{i})$ term in place of $\mu$, the slight difference in definition between $\operatorname{ampl}_{k}(\hat{z}_{i})$ and $\Gamma_{k}$, and the extra $1+2^{-p_{i}}$ term.
-These differences are small but necessary for soundness.
-[^11] In fact, the name $\operatorname{ampl}$ is a pun on both “amplification factor” and ”ample” as in sound.
+This in turn causes small, but necessary, [^10] In fact, the name $\operatorname{ampl}$ is a pun on both “amplification factor” and ”ample” as in sound.
+differences like the $\operatorname{intro}(\hat{z}_{i})$ term in place of $\mu$, the slight difference in definition between $\operatorname{ampl}_{k}(\hat{z}_{i})$ and $\Gamma_{k}$, and the extra $1+2^{-p_{i}}$ term.
 4.3. Precision Tuning
@@
 To actually use Equation 2 we must compute $\operatorname{intro}(\hat{z}_{i})$ and $\operatorname{ampl}_{k}(\hat{z}_{i})$, which requires optimizing over $\varepsilon^{*}$.
-However, we can use interval arithmetic to perform this optimization.
+We can use interval arithmetic to perform this optimization.
 Moreover, already-computed interval outputs from each operation can be reused to make this fast.
@@
 So,
-$\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{f(x_{i}^{*},y_{i}^{*})}{f(x_{i},y_{i})}\right|\right\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{z_{i}^{*}}{z_{i}}\right|\right\rceil\in\lceil\log_{2}|\bar{z}_{i}/\bar{z}_{i}|\rceil$
+$\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{f(x_{i}^{*},y_{i}^{*})}{f(x_{i},y_{i})}\right|\right\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{z_{i}^{*}}{z_{i}}\right|\right\rceil\in\lceil\log_{2}|\bar{z}_{i}/\bar{z}_{i}|\rceil.$
 Note that $\bar{z}_{i}/\bar{z}_{i}$ is an interval-arithmetic division, so it isn’t equal to 1; it’s an interval that contains 1, but also some larger and smaller values.
@@
 In other words, $\operatorname{intro}$—and, with a similar trick, $\operatorname{ampl}_{k}$—can be computed with standard interval arithmetic.
-The upper bound on $\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil$ is the high endpoint of $\lceil\log_{2}|\bar{z}_{i}/\bar{z}_{i}|\rceil$.
 Write $\operatorname{lo}(\bar{x})$ and $\operatorname{hi}(\bar{x})$ for the high and low endpoint of an interval $\bar{x}$, and assume for now (see Section 6) that $\bar{x}$ contains neither zero nor infinity.
@@
 We call this the “exponent trick”: bounding the logarithms of $\operatorname{intro}$ and $\operatorname{ampl}_{k}$ using only: 1) integer operations, 2) on $\operatorname{minlog}$ and $\operatorname{maxlog}$, 3) of already-computed intervals $\bar{x}$, $\bar{y}$, and $\bar{z}$.
-The remainder of this section describes exponent tricks for various operators.
+The remainder of this section describes exponent tricks for various operators’ $\operatorname{ampl}_{k}$ terms.
 Some of the derivations are mathematically challenging, but all the key results are shown in Tables 2 and 2.
@@
 Division, square and cube roots, exponents, and logarithms allow a similar algebraic rearrangement, producing a formula that uses only integer operations on $\operatorname{maxlog}$ and $\operatorname{minlog}$ of $\bar{x}_{i}$, $\bar{y}_{i}$, and $\bar{z}_{i}$.
-The resulting formulas are shown in Table 2.
+The resulting “exponent tricks” formulas are shown in Table 2.
 5.3. Approximations
-However, for some operators, an approximation has to be performed to bound $(\partial_{k}f)(x,y)$ via some expression over $x$, $y$, or $z$.
-This approximation allows us to construct an exponent tricks formula for $\operatorname{ampl}_{k}(f)$, that is, a formula that bounds $\lceil\log_{2}\operatorname{ampl}_{k}(f)\rceil$ using only integer operations on $\operatorname{maxlog}$ and $\operatorname{minlog}$ of $\bar{x}$, $\bar{y}$, and $\bar{z}$.
+However, for some operators, an approximation has to be performed to bound $(\partial_{k}f)(x,y)$ via some expressions over $x$, $y$, or $z$.
+These approximations allow us to construct exponent tricks formulas for $\operatorname{ampl}_{k}(f)$, that is, formulas that bound $\lceil\log_{2}\operatorname{ampl}_{k}(f)\rceil$ using only integer operations on $\operatorname{maxlog}$ and $\operatorname{minlog}$ of $\bar{x}$, $\bar{y}$, and $\bar{z}$.
 Table 2 shows all of the approximations used and the exponent trick formulas they result in, and readers are invited to focus their attention on it.
@@
 The amplification factor for $\sin(x)$ involves $\cos(x)$. Because $|\cos(x)|\leq 1$, we have $\log_{2}|\cos(x^{*})|\leq 0$; this allows us to show that
-$\lceil\log_{2}\operatorname{ampl}_{1}(\sin(\hat{x}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{y}.$
+$\lceil\log_{2}\operatorname{ampl}_{1}(\sin(\hat{x}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{z}.$
 Likewise, for cosine we must bound $\sin(x)$; one simple approximation is $|\sin(x)|\leq\min(|x|,1)$, meaning $\log_{2}|\sin(x^{*})|\leq\min(\operatorname{maxlog}(\bar{x}),0)$.
 This produces the approximation
-$\lceil\log_{2}\operatorname{ampl}_{1}(\cos(\hat{z}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{y}+\min(\operatorname{maxlog}(\bar{x}),0).$
-These approximations are sound and easy to compute. Moreover, the approximation error is small when either the input or the output is near zero, which helps reduce the needed precision.
+$\lceil\log_{2}\operatorname{ampl}_{1}(\cos(\hat{x}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{z}+\min(\operatorname{maxlog}(\bar{x}),0).$
 Power
-The power function $x^{y}$ has an amplification factor for $y$ that uses a $\log(x^{*})$ term. This means that we must bound $\log_{2}|\log(x)|$ in terms of $\log_{2}|x|$. To do so, we first rewrite
+The $\operatorname{ampl}_{2}$ factor for the power function $x^{y}$ has uses a $\log(x^{*})$ term.
+This means that we must bound $\log_{2}|\log(x)|$ in terms of $\log_{2}|x|$.
+To do so, we first rewrite
 $\log_{2}|\log(x)|=\log_{2}\left|\frac{\log_{2}(x)}{\log_{2}(e)}\right|=\log_{2}|\log_{2}(x)|-\log_{2}\log_{2}e\leq\log_{2}|\log_{2}(x)|-\frac{1}{2}.$
@@
 Hyperbolic Functions
-The amplification factors for $\sinh$ and $\cosh$ both require bounding $\tanh(x)$.
-In both cases, we use the well-known fact that $|\tanh(x)|\leq\min(|x|,1)$, leading to $\lceil\log_{2}|\tanh(x)|\rceil\leq\min(\operatorname{maxlog}(\bar{x}),0)$.
+The amplification factors for $\sinh$ and $\cosh$ both use $\tanh(x)$.
+Because $|\tanh(x)|\leq\min(|x|,1)$, we have $\lceil\log_{2}|\tanh(x)|\rceil\leq\min(\operatorname{maxlog}(\bar{x}),0)$.
 This approximation is relatively tight, producing at most one bit of error.
-The amplification factor for $\tanh(x)$ requires bounding $x/\sinh(x)\cosh(x)$.
-We use a very weak bound, $x/\sinh(x)\cosh(x)\leq 1$.
-This over-approximates significantly for large values of $x$, though not near 0.
+The amplification factor for $\tanh(x)$ requires bounding $x/\sinh(x)\cosh(x)$. We use a very weak bound, $x/\sinh(x)\cosh(x)\leq 1$, which is only tight near 0.
 Tangent
@@
 $=\operatorname{maxlog}\bar{x}_{i}+|\operatorname{maxlog}\bar{z}_{i}|+\operatorname{logspan}\bar{z}_{i}+1,$
-which now only uses already-computed intervals. Moreover, this approximation is off by at most one bit.
+which now only uses already-computed intervals and is at most one bit from being tight.
 Arctangent
@@
 Table 1. Bounds on amplification factors using only the exponents of already-computed intervals $\bar{x}$, $\bar{y}$, and $\bar{z}$. The formulas in this table do not use any approximations.
-Operator ($f$) Approximation used & exponent tricks bound
+Operator ($f$) Approximation used & exponent tricks bound for $\lceil\log_{2}\operatorname{ampl}_{k}\rceil$
 $x^{y}$, $k=2$ $\log_{2}|x|\leq|x|-\frac{1}{2}$
@@
 […]
-$\operatorname{maxlog}(\bar{x})+\operatorname{maxlog}(\bar{y})-2\max(\operatorname{minlog}(\bar{x}),\operatorname{minlog}(\bar{y}))-\operatorname{minlog}(\bar{z})$
+$\operatorname{maxlog}(\bar{x})+\operatorname{maxlog}(\bar{y})-2\min(\operatorname{minlog}(\bar{x}),\operatorname{minlog}(\bar{y}))-\operatorname{minlog}(\bar{z})$
 Table 2.
@@
 The bounds for $\sin$ and $\cos$ over-approximate by more, though on relatively rare inputs.
-6. Implementation Details
-Reval provides a simple user interface: the user first compiles an expression, which may have free variables, and then applies it to specific values.
-The compilation produces a simple register machine with a vector of instructions and separate vectors with, for each instruction, an output interval and a precision assignment.
+6. Implementing Real Evaluation
+Reval provides a simple user interface: the user first compiles (one or more) expressions, which may have free variables, and then applies it to specific values for those free variables.
+The compilation produces a simple register machine with a vectors of instructions, their precision assignments, and their outputs.
 6.1. Compilation
-Reval supports expressions with real-number operations, and real constants, boolean operations, and if statements, but no loops.
-Specifically, Reval supports the constants $\pi$ and $e$, the standard arithmetic operators (plus, minus, times, divide), and every function specified in the POSIX standard for math.h, including square and cube roots; exponents, powers, and logarithms; trigonometric, arc-trigonometric, and hyperbolic and arc-hyperbolic functions; the error and gamma functions; modulus and remainder operators; rounding functions; and miscellaneous others like $\min$ and $\max$.
-It also supports comparison functions and boolean connectives.
-Thanks to the lack of loops, each expression can be flattened to a linear instruction sequence.
-Users can also supply a batch of multiple expressions to evaluate on the same inputs.
-Reval performs common subexpression elimination so this can dramatically reduce runtime if the expressions have shared structure, though this feature is not used in this paper.
-Instructions are written in two-argument form, reading in their arguments from input registers and returning a result that is saved in the instruction’s output register.
-Interval operators come from the Rival library (Flatt and Panchekha, 2023), which has a couple of unusual features.
-It provides boolean functions like comparisons and connectives that output ternary “boolean intervals”; for example, $[1,1]<[0,2]$ returns the “maybe” interval [,]bottomtop[$\bot$,$\top$][ ⊥ , ⊤ ].
-An if statement with a “maybe” condition unions the intervals in both branches.
-In the presence of domain errors, Rival intervals enclose all outputs on valid inputs; for example, $\sqrt{[-1,1]}=[0,1]$.
-A special empty interval $\emptyset$ is used when the whole input interval is invalid.
-All intervals also track domain errors using a ternary error flag; for $\sqrt{[-1,-1]}$ the error flag is $\top$ and for $\sqrt{[-1,1]}$ the error flag is [,]bottomtop[$\bot$,$\top$][ ⊥ , ⊤ ].
-Rival also provides movability flags (Flatt and Panchekha, 2023) that detect some unevaluable inputs.
-Reval typically attempts to correctly round each expression to double precision.
-However, this rounding behavior is customizable by supplying a discretization.
-A discretization is a type $T$, a target precision $t$, a conversion function from MPFR values to $T$s, and a distance function on pairs of $T$s.
-The distance function is called with the two converted $T$ endpoints of an interval and returns one of three values: the endpoints can be equal, unequal, or neighbors, meaning that the two $T$ values share a rounding boundary.
-The “equal” return value is used for correct rounding; if the user instead wants faithful rounding, they can supply a distance function that returns “neighbor” instead of “equal”.
+Reval supports expressions containing the constants $\pi$ and $e$, the standard arithmetic operators (plus, minus, times, divide), and every function specified in the POSIX standard for math.h, including square and cube roots; exponents, powers, and logarithms; trigonometric, arc-trigonometric, and hyperbolic and arc-hyperbolic functions; the error and gamma functions; modulus and remainder operators; rounding functions; and miscellaneous others like $\min$ and $\max$.
+It also supports comparison functions, boolean connectives, and if statements.
+Interval operators are provided by the Rival library (Flatt and Panchekha, 2023), which also provides error flags for detecting domain errors and movability flags for detecting unevaluable expressions.
+Reval compiles the expression(s) to a register machine instruction sequence in three-argument form and performs common subexpression elimination.
+By default, Reval evaluates each expression to correctly-rounded double precision values.
+This rounding behavior is customizable by supplying a discretization: a type $T$, a target precision $t$, a conversion from MPFR to $T$s, and a distance function on pairs of $T$s.
+The distance function is called with the two converted $T$ endpoints of an interval and decides whether the endpoints are equal, unequal, or neighbors, where the last is used for correct rounding.
 6.2. Evaluation
-Reval evaluates each instruction in turn, at that instruction’s precision assignment.
-Then, for each output register, Reval uses the conversion and distance functions to determine if the computed intervals are sufficiently narrow.
-If not, Reval performs a precision assignment using the exponent trick formulas and the already-computed intervals for each register, and repeats this process until a sufficiently-narrow interval is found.
+Each instruction in the instruction sequence has its own precision, initialized to a constant.
+To evaluate the compiled expression, Reval repeatedly evaluates each instruction and then tunes the precision assignment, stopping once each output register holds an interval with equal endpoints under the discretization’s distance function.
 This minimal evaluation loop reduces evaluation overhead, and performance is largely determined by the choice of precision for each instruction.
-If the precision assigned to a specific instruction does not change (or even decreases) from one iteration to the next, there’s no need to recompute that instruction, as long as its inputs are not recomputed.
-Reval checks for this condition during precision assignment and sets a bit in a per-instruction “repeat vector”; instructions with the repeat bit set are not re-executed.
-On the other hand, if an operation is ever assigned a precision higher than a user-specified precision bound, Reval terminates execution and signals an error.
-This “early stopping” can trigger as early as the first iteration and saves significant time on points Reval cannot evaluate.
-Reval can also exit the evaluation loop if the output intervals’ error flag is set to $\top$, if movability flags prove the expression unevaluable, or if a user-defined iteration limit is reached.
-6.3. Precision Tuning
 Reval’s precision tuning pass is basically Equation 3 converted to code.
-Reval allocates a vector $P[z_{i}]$ and fills the vector starting from the end.
+Instruction precisions are stored in a vector $P[z_{i}]$ which is filled in a “backward pass” starting from the end.
 The already-filled, later elements of the array represent the precisions $p_{i}$ for each instruction, while the initial, un-filled elements represent a target precision for $\operatorname{bound}(\hat{z}_{i})$.
@@
 The $\operatorname{intro}$ and $\operatorname{ampl}$ terms are computed using the exponent trick formulas of Tables 2 and 2.
-This single “backward pass” over the precision array suffices to compute a precision assignment for each instruction.
 Reval performs some additional, minor optimizations to make this formula even faster to compute, including pre-propagating the “${}+2$” factors and computing the $\operatorname{intro}$ and $\operatorname{ampl}$ terms simultaneously.
-Soundness ensures that Reval terminates: without a sound precision assignment, it would be possible to enter an infinite loop, where at each iteration the same precisions are assigned.
-Reval does include code to check for this condition (in case the user provides a custom discretization with a too-low target precision $t$) but this code is not triggered in our evaluation.
-However, achieving termination also requires some additional pragmatic changes to the simple theory of Equation 3.
-6.4. Slack
-Sometimes, an interval crosses zero, in which case $\operatorname{minlog}(\bar{z})$ is not defined—it approaches negative infinity.
-Similarly, some intervals are unbounded, in which case $\operatorname{maxlog}(\bar{z})$ approaches positive infinity.
-In both cases, the exponent trick formulas would evaluate $\operatorname{intro}$ and $\operatorname{ampl}_{k}$ as infinity, which of course is not a usable precision assignment.
+If the precision assigned to a specific instruction does not change (or even decreases) from one iteration to the next, there’s no need to recompute that instruction, as long as its inputs are not recomputed.
+Reval checks for this condition during precision assignment and sets an instruction-specific bit in a “repeat vector”; instructions with the repeat bit set are not re-executed.
+This “repeat optimization” is only possible thanks to the non-uniform precisions; with a uniform precision, every instruction’s precision would increase at every iteration and no repeat bits would be set.
+6.3. Slack and Correct Rounding
+Since Reval’s precision assignment algorithm is sound, the precisions assigned in the precision tuning pass are high enough that the next iteration will yield narrow-enough intervals produce a faithfully-rounded result.
+However, if an intermediate interval $\bar{z}$ either crosses zero or contains infinity, its $\operatorname{minlog}(\bar{z})$ or $\operatorname{maxlog}(\bar{z})$ approach infinity, which makes precision assignment impossible.
 Intuitively, what’s happening is that the precision assignment algorithm does not know how many bits will suffice to eliminate, say, a catastrophic cancellation, and assigns infinitely many bits to remain sound.
-In this case, Reval falls back to Ziv onion peeling; this is why Reval can require multiple iterations of precision assignment and evaluation before finding a narrow-enough interval.
-Specifically, Reval uses $\operatorname{slack}[n]$ in place of any infinity in the exponent trick formulas, where $n$ is how many iterations Reval has performed and $\operatorname{slack}[n]$ is a guess as to how many bits might be needed.
+Reval addresses this issue with a partial return to precision doubling, replacing the infinite value with a guess.
+Specifically, Reval uses $\operatorname{slack}[n]$ in place of any infinity in the exponent trick formulas, where $n$ is how many iterations Reval has performed.
 So, for example, the interval $\bar{z}=[-2^{-93},2^{-108}]$ has $\operatorname{minlog}(\bar{z})=\min(-93,-108)+\operatorname{slack}[n]$.
-Slack is equal to 512 after the first iteration, and doubles for every further iteration.
-This slack mechanism reintroduces a naive precision-doubling algorithm, but only partially:
-  • The precisions are still not uniform;
-  • The non-zero, non-infinite endpoints still provide useful information about the necessary precision, so even though slack terms involve guesses they are at least somewhat informed; and
-  • Slack typically affects only one or a few operations, and the other operations still set their precisions using sound mechanisms.
-Because slack only affects a few operations, it can be set more aggressively. The initial guess of 512, for example, is much larger than Sollya’s initial guess of 64.
-6.5. Correct rounding
-Reval also uses the slack mechanism to achieve correct rounding.
-Correct rounding requires firmly placing the value to one side or another of a rounding boundary in the target precision.
-The challenge is that the output can be very close to such a rounding boundary, so that proving it to be on one side or another can require $\bar{z}$ to be much narrower than merely $2^{-p}$.
-Note that this is rare: in most cases, the two endpoints are either more than 1 ULP apart or match exactly; however, correct rounding is valuable for a variety of use cases (Brisebarre et al., 2024) due to its determinism.
-Reval uses the slack mechanism to handle this case: if the output interval rounds to two successive floating-point numbers in the target precision, meaning that Reval has achieved a faithfully- but not correctly-rounded result, the target precision $t$ is increased by $\operatorname{slack}[n]$ before performing precision assignment.
+While this “slack mechanism” is similar in some ways to the standard precision-doubling algorithm, it has some key advantages.
+Precisions are still not uniform, and the non-zero, non-infinite endpoints of intervals still provide useful information to precision tuning.
+Moreover, slack typically affects only one or a few operations, and all other operations still set their precisions using sound mechanisms.
+Because of this last point, it can be set more aggressively: 512 for $\operatorname{slack}[0]$ and doubling in later iterations.
+Even a precision assignment with no slack still only guarantees a faithful rounding, not a correct rounding, because even a very narrow output interval can still straddle a rounding boundary.
+Reval detects this case using the “neighbor” distance value and increases the target precision until the output interval falls to one side of the rounding boundary or the other.
+Since there’s no principled way to know how close to the rounding boundary the true value lies, Reval increases the target precision by $\operatorname{slack}[n]$.
+Users can disable this mechanism by providing a custom discretization which never returns the “neighbor” distance value, but it is on by default because correct rounding is valuable for a variety of use cases (Brisebarre et al., 2024).
+6.4. Invalid Inputs
+Inputs to a real expressions can cause domain errors (detected by Rival’s error flags) or simply be impossible to evaluate under a given precision bound. Reval thus exits the evaluation loop if:
+  • an output register’s error flag is set;
+  • an output register’s movability flags prove the expression unevaluable at any precision;
+  • an operation is assigned a precision beyond a user-specified precision bound. [^13] Reval also supports a user-set iteration bound.
+The last case, “early exit”, can be triggered as early as the first tuning iteration and is critical for Reval’s robustness to invalid inputs.
+That said, it can trigger inappropriately: the precision assignment over-estimates the necessary precision, so a lower precision, possibly lower than the precision bound, might still work.
+This edge case is rare but does occur in our evaluation.
+To help address this issue, Reval has an optional “lower bound mode”.
+In this mode, a second set of exponent tricks formulas, which under-approximate the amplification factors, are used for the early exit calculation.
+This second set of formulas largely with switches $\operatorname{maxlog}$ for $\operatorname{minlog}$ and vice versa, with adjustments to the approximations used for some operations.
+If the precision assignment is larger than the precision bound but the lower-bound formula is below the precision bound, the precision assignment is set to exactly the precision bound.
+This avoids all cases of inappropriate early stopping in our evaluation, but also increases the time spent on unevaluable inputs, so this “lower bound mode” is off by default.
 6.5. Other pragmatic tweaks
-Rival, the interval library Reval uses, does not seem to provide operators that are tight in the sense of IEEE 1788 (iee, 2015).
-This means that Equation 3 can produces slightly-too-low precision assignments.
-In practice, this is not much of a problem, because the interval operators that aren’t tight, like $\sin$ and $\cos$, also use over-approximating exponent trick formulas.
-But as extra security to avoid issues, Reval uses “${}+5$” instead of “${}+2$” in its precision assignment formulas.
+Equation 3 assumes the interval library produces maximally tight intervals (iee, 2015), but this is not actually true for Rival.
+As extra security to avoid issues, Reval thus uses “${}+5$” instead of “${}+2$” in its precision assignment formulas.
 This allows for up to 3 ULPs of looseness in each operator, which we believe is more than Rival needs.
-[^14] We’ve validated this with some experiments, but to our knowledge Rival has not been proven correct and verified bounds are not available.
+[^14] We’ve validated this empirically but Rival has not been proven correct and verified bounds are not available.
 Since much of runtime is spent at very high precisions, the addition of a few bits doesn’t impact performance much.
@@
 For these operators Reval simply uses $\operatorname{slack}$ for their $\operatorname{ampl}$ factors.
-Naturally, Reval’s performance is then no better than naive precision-doubling, but these operators are rare so basic support seems acceptable.
-For some operators, the slack mechanism is only resorted to for some inputs.
-For example, $\lceil\log_{2}\operatorname{ampl}_{1}(\operatorname{asin}(\hat{x}_{i}))\rceil\leq\operatorname{maxlog}(\bar{x}_{i})-\operatorname{minlog}(\bar{z}_{i})+1$ as long as $\operatorname{maxlog}(\bar{z}_{i})\leq 0$, but when that condition is not true, slack still has to be added.
-Similarly, for the rounding operators like $z=\lfloor x\rfloor$, $\lceil\log_{2}\operatorname{ampl}_{z}(\hat{z})\rceil$ is set to $\operatorname{maxlog}(\bar{x})$ if $\hat{z}$ contains more than one integer, $\infty$- ∞ if it contains zero integers, and $\operatorname{slack}[n]$ if it contains exactly one integer (meaning the true value of $x$ might be right near a rounding boundary).
-Finally, the $\operatorname{logspan}$ terms in the exponent trick formulas caused issues in early iterations when intervals are very wide.
-We therefore set $\operatorname{logspan}(\bar{z})=0$; surprisingly, this does not affect soundness, because in the final iteration, where the slack mechanism isn’t used, all values must be known to at least 2 bits of precision, meaning $\operatorname{logspan}(\bar{z})\leq\frac{1}{4}$.
-Since no exponent trick formula contains more than three $\operatorname{logspan}$ terms, the total added precision from $\operatorname{logspan}$ terms is less then 1.
-Reval’s use of “${}+5$” instead of “${}+2$” covers this extra bit as well.
+Reval is then no better than standard precision-double algorithms, but these operators are rare so basic support seems acceptable.
+Alternatively, some operators like $\operatorname{asin}$ use the slack mechanism only for certain inputs.
+That is, $\lceil\log_{2}\operatorname{ampl}_{1}(\operatorname{asin}(\hat{x}_{i}))\rceil\leq\operatorname{maxlog}(\bar{x}_{i})-\operatorname{minlog}(\bar{z}_{i})+1$ when $\operatorname{maxlog}(\bar{z}_{i})\leq 0$, but when that condition is not true, slack is used.
+Similarly, for the floor operator, $\lceil\log_{2}\operatorname{ampl}_{z}(\lfloor\hat{x}\rfloor)\rceil=\operatorname{maxlog}(\bar{x})$ if $\hat{x}$ contains more than one integer, $\infty$- ∞ if it contains zero integers, and $\operatorname{slack}[n]$ if it contains exactly one integer.
+Finally, the $\operatorname{logspan}$ terms in the exponent trick formulas are very large in early iterations; we therefore replace $\operatorname{logspan}(\bar{z})\mapsto 0$.
+Surprisingly, this does not affect soundness, because in the final iteration with no slack mechanism, all values are known to at least 2 bits of precision, meaning $\operatorname{logspan}(\bar{z})\leq\frac{1}{4}$.
+Since no exponent trick formula contains more than three $\operatorname{logspan}$ terms, the total added precision from $\operatorname{logspan}$ terms is less then 1, which is covered by Reval’s use of “${}+5$” instead of “${}+2$”.
 7. Evaluation
@@
   1. RQ1
-    Can Reval evaluate mathematical expressions faster than a uniform-precision baseline?
+    Can Reval evaluate expressions faster than a uniform-precision baseline?
   2. RQ2
-    Can Reval evaluate more mathematical expressions than a uniform-precision baseline?
+    Is this because Reval uses lower precisions than a uniform-precision baseline?
   3. RQ3
-    Does Reval use lower precisions, on average, than a uniform-precision baseline?
-We evaluate all three research questions on a set of 528 benchmarks drawn from the Herbie 2.0 benchmark suite, each evaluated at 8256 inputs.
-For the uniform-precision baseline, we use both Reval with precision tuning disabled and also the state-of-the-art Sollya tool (except for RQ3).151515 Other tools that perform real evaluation include Wolfram Mathematica (Wolfram, 2020), $\lambda_{s}$ (Sherman et al., 2021), the Android and Windows calculators (Boehm, 2020), Calcium (Johansson, [n.
-d.]), and possibly others.
-Prior work (Flatt and Panchekha, 2023) has found Mathematica to be much slower than Reval, and the other libraries are not, to our knowledge, optimized for performance.
+    How do Reval’s precision assignments contribute to performance improvements?
+We evaluate all three research questions by comparing Reval to a variant, Baseline, that uses a standard uniform precision-doubling algorithm, on 567,238 inputs to 469 benchmarks from the Herbie 2.0 benchmark suite.
+For RQ1, we also compare to the state-of-the-art Sollya tool.151515 Wolfram Mathematica (Wolfram, 2020), the $\lambda_{s}$ (Sherman et al., 2021) and Marshall languages (Sherman et al., 2019), the CGAL and LEDA graphics packages (Seel, 2024; Näher and Mehlhorn, 1992), the Calcium (Johansson, [n.
+d.]) library, and the Android and Windows calculators (Boehm, 2020) also perform real evaluation but aren’t optimized for performance and use the similar uniform precision algorithms to Sollya and Baseline.
 7.1. Methodology
-The 528 benchmarks are mathematical expressions written in the FPCore 2.0 format (Damouche et al., 2016).
+Benchmarks
+The Herbie 2.0 suite contains 509 mathematical expressions written in the FPCore 2.0 format (Damouche et al., 2016).
 Benchmarks are short, with 1–59 operations (average 11.6) over 1–9 variables (average 2.5).
-The most common operations by far are the arithmetic operations, with multiplication at over 30% of operations and addition, subtraction, and division constituting another 50% of operations.
-The most common non-arithmetic operations are the exponential, power, sine, cosine, logarithm, and square root operations, each approximately 1.5% of operations.
-The rarest operations were arctangent, hypotenuse, and fused multiply-add.
-For each benchmark, we randomly sample 8256 input points (that is, values for each variable) using the Herbie 2.0 tool and use the same input points for Reval and both baselines.
-Input points are sampled using rejection sampling, with inputs that have a domain error according to the baselines or Reval being rejected.
-We discard benchmarks with loops, conditionals, data structures (arrays or tensors), or functions not supported by Sollya (mainly cbrt and atan2), which eliminates 17 of the full 547-strong benchmark suite.
-In addition, two benchmarks (that we know of) trigger a bug in Sollya related to the subexpressions $z^{z}e^{-z}$ and $e^{-z}\ell^{\exp(z)}$, reported to and acknowledged by the developers.
-We remove these two benchmarks as well.
-528 remain, of which 29 are evaluated to single-precision accuracy, while 499 are evaluated to double-precision accuracy; inputs are also given in the same precision.
-Sollya doesn’t provide hypot and fma functions so these are expanded into $\sqrt{x^{2}+y^{2}}$ and $x\cdot y+z$ during compilation to Sollya’s input language.
-98 benchmarks specify a precondition, some of which are simple ranges but others of which involve complex conditionals.
-Since Sollya does not support preconditions, we do not time precondition evaluation, but use only input points that satisfy the preconditions.
-Reval is set to correctly round each expression on each of its inputs; however, Sollya instead only attempts to round faithfully.161616 Sollya’s definition of faithful rounding allows it to round a value to any of three possible floating-point values, presumably in the case of exact floating-point results.
-This is rare but does happen in our evaluation.
-Moreover, in some cases, the correctly-rounded result is infinite (for example, if the real result is $10^{400}$).
-In this case, Reval returns infinity but Sollya’s results vary; we track these input points separately and ignore them in RQ1–3.
+The most common operations by far are the arithmetic operations (80%+) but the exponential, power, sine, cosine, logarithm, and square root operations are also common (each $\approx 1.5\%$).
+We discard 14 benchmarks with unsupported operations [^16] Loops, conditionals, data structures (arrays or tensors), and the cbrt and atan2 functions not supported by Sollya.
+as well as two benchmarks that trigger a bug in Sollya, [^17] Sollya incorrectly handles underflow $z^{z}e^{-z}$ and $e^{-z}\ell^{\exp(z)}$; the bug was reported and acknowledged by the developers.
+leaving 493 total.
+Inputs
+For each benchmark, we randomly sample 8256 double-precision input points (that is, values for each variable) using the Herbie 2.0 tool [^18] Herbie 2.0 uses rejection sampling and samples points with no domain errors and which satisfy any benchmark-specific preconditions.
+and use the same input points for Reval and the baselines.
+All inputs to the remaining 493 benchmarks are evaluated to correctly-rounded double-precision values;191919 Sollya mostly produces correctly-rounded results, but in 5355 cases, roughly 1%, instead produces a faithful rounding.
+inputs whose correctly-rounded result is infinite are discarded because Sollya’s handling of these points is inconsistent.
+Of the remaining 4,070,208 input points, the vast majority can be evaluated at the default starting precision and do not require precision tuning; we discard these points as well, leaving a total of 567,238 inputs across 469 benchmarks.
 Baselines
-All tools are run with Sollya’s default minimum and maximum precisions or $10+t$ and $(10+t)\cdot 512$, where $t$ is either 24 or 53.
-These initial and maximum precisions are Sollya’s defaults.
-For Sollya the necessary precision indicated by setting the prec variable.
-Sollya’s main evaluation strategy is to evaluate the whole expression under a uniform precision with interval arithmetic, doubling the precision until a faithful rounding is found.
-However, Sollya does, in various special cases, apply additional tricks, including (from an examination of the source code) Taylor series, l’Hôpital’s rule, and in some cases smaller precision changes than doubling.
-To the best of our knowledge, however, it does not perform mixed-precision evaluation, except in the case of constants that are cached at high precision.
-Reval applies the precision tuning algorithm described in this paper, with failure after 5 iterations.
-We also compare to “Baseline”, which is Reval but applying a uniform precision that doubles with each iteration until the correct rounding has been obtained or the maximum precision has been reached.
-For each tool, benchmarks are compiled from FPCore to the tool’s input syntax.
-The compilations are all straightforward.
-For Sollya each expression is first defined as a function, with the compilation time not measured, and then that function is evaluated on each input point in turn.
-The separate compilation step allows Sollya to perform compile-time optimization or to share caches across inputs.
-Reval also performs a separate compilation step; however, it does not perform any caching except for exact constants like 1.
-To check that the compilation is performed correctly, we compare the outputs from all three tools (which is how we discovered the previously-mentioned Sollya bug).
+All 567,238 are evaluated using both Reval as well as Baseline, a variant of Reval which uses a uniform precision that doubles iteration to iteration.
+Since Reval and Baseline are the same codebase, just with different tuning strategies, this allows a precision evaluation of Reval’s precision-tuning strategy.
+To ensure Reval and Baseline are competitive implementations, we also compare both against Sollya, an independently-developed environment for floating-point code development which also uses a uniform precision.202020 Sollya typically doubles the precision in each iteration, but in some cases can detect when it’s “close” to the correct precision and increase the precision by less than doubling.
+In any case, the precision is uniform.
+Reval and Baseline consume the benchmark’s standard FPCore formats, but for Sollya we also compile each benchmark into Sollya’s input format; we check that all three tools agree on each input to ensure that the compilation is correct.
 Measurement
-Evaluation time for an input point must be measured carefully.
-Reval and Baseline are called within the same process, so a simple nanosecond timer is used.
-Sollya, however, is started as an external process.
-To avoid including parsing time and other overhead, Sollya’s time command is used to measure its running time.
-All tools are subjected to a 20 millisecond timeout.
-Sollya in particular suffers from what seem like infinite loops; 20 milliseconds was selected as the shortest timeout that still allows Sollya to successfully evaluate most of even the most difficult inputs.
-A larger timeout would only bias the results toward Reval.
+For each tool we measure evaluation time and compilation time separately, and focus on evaluation time.
+Reval and Baseline are called within the same process, so a simple nanosecond timer is used, while for Sollya we use its time command.
+All three tools are subject to a 20 millisecond timeout per input points.
+A short timeout is chosen to avoid biasing results due to rare seemingly-infinite loops in Sollya; the timeout is not significant for Reval or Baseline.
 Evaluation is performed on a machine with an Intel i7-8700 at 3.7 GHz with 32 GB of DDR4 memory running: Ubuntu 22.04; Sollya 7.0 with MPFI 1.5.3, MPFR 4.1.0, and GMP 6.2.1; and Reval 2.0 with Racket 8.11.1.
@@
 RQ1
-Figure 3 shows our results for RQ1.
-The figure shows evaluations per second for Reval and Baseline, relative to Sollya.
-Only the 4,359,168 input points where all three tools successfully evaluate the expression are included.171717 This way the handling of invalid inputs, unsamplable inputs, precision limits, and so on do not influence the results.
-Input points are aggregated by “difficulty”, which is the number of iterations Reval needs to evaluate the point; higher difficulty typically means higher precision.
-Note that this difficulty metric (chosen because it is easy to compute) is biased against Reval: a point that takes Reval many iterations may nonetheless be easy for Sollya or Baseline.
-Very roughly, every difficulty level has about 20×20$\times$20 × fewer points than the previous; there are only 30 difficulty-5 points, so its timing results are somewhat unreliable, the the relative standings of the three tools occur repeatedly.
-Reval achieves significantly more evaluations per second, both on average and at any difficulty level.
-Averaged over all points, Reval is 1.25×1.25$\times$1.25 × faster than Sollya and 1.21×1.21$\times$1.21 × faster than Baseline.
-In fact, Baseline and Sollya perform at similar speeds across the difficulty spectrum, as expected given their similar precision assignment strategy.
-Reval has the smallest advantage at difficulty 2, which we believe is due to the slack mechanism sometimes overestimating the necessary precision.
-Reval’s advantage grows as difficulty increases: for the most difficult points, Reval is 2.99×2.$9$ faster than Sollya and 1.71×1.71$\times$1.71 × faster than Baseline.
-This is expected: more difficult points require higher precisions meaning that precision assignment provides a greater benefit.
+Figure 3 shows evaluations per second for Reval, Baseline, and Sollya, aggregated by Baseline’s chosen precision212121 This is typically also the uniform precision used by Sollya, and typically the largest precision assigned by Reval.
+and normalized against Baseline’s runtime.
+Reval is significantly faster on average, by 1.62×1.62$\times$1.62 × compared to Baseline and by 1.47×1.47$\times$1.47 × compared to Sollya.
+Inputs evaluated at higher-precision are typically harder; Reval is faster than Baseline at any precision and faster than Sollya past $2^{9}$ bits of uniform precision, with Reval’s advantage growing with precision: for the most precision-demanding inputs, Reval is 4.92×4.92$\times$4.92 × faster than Sollya and 2.38×2.38$\times$2.38 × faster than Baseline.
+In general, Sollya is somewhat faster than Baseline (though note the discussion on timeouts); we suspect this is due to Sollya’s more-careful memory management, orthogonal to Reval’s precision assignment algorithm.
 Figure 2.
-Reval’s, Sollya’s, and Baseline’s speed by difficulty of input point.
-Sollya is normalized to 1.0; higher is better.
-All three tools perform similarly on the easiest points, and Baseline and Sollya perform similarly throughout, as expected given their very similar core algorithms.
-However, Reval has a growing advantage on harder inputs, with a speed-up over Sollya of 1.25×1.25$\times$1.25 × overall and 2.99×2.$9$ on the most difficult inputs.
+Evaluations per second, with Baseline normalized to 1.0; higher is better.
+Reval is, on average, 1.62×1.62$\times$1.62 × faster than Baseline and 1.47×1.47$\times$1.47 × faster than Sollya.
+Inputs are aggregated by Baseline precision, with higher precisions indicating “harder” inputs; Reval has a growing advantage, up to 2.38×2.38$\times$2.38 × over Baseline and 4.92×4.92$\times$4.92 × over Sollya.
 Figure 3.
-The number of points evaluated by Reval, Sollya, and Baseline, broken down by whether the correctly-rounded result was zero or a non-zero real number.
-Higher is better.
-Reval and Baseline both evaluate far more points than Sollya.
-Reval and Baseline are largely comparable, showing that Reval’s precision tuning does not much affect whether points are successfully evaluated.
+A weighted histogram showing the distribution of per-operation execution time by precision, for both Reval and Baseline; lower is better.
+Baseline wastes a lot of time re-computing the same operations at successively higher precisions, leading to a lot of wasted time at precisions between $2^{6}$ and $2^{11}$.
+The “tuning” column shows the run time of Reval’s precision tuning algorithm, over all inputs and iterations.
 RQ2
-Figure 3 shows our results for RQ2.
-This graph considers only points successfully evaluated by at least one, but not all, of the tools.
-Reval and Baseline evaluate a similar number, with Reval evaluating more points where the correct answer is 0 (3743 to Baseline’s 3741) but fewer points where the correct answer is non-zero (10 to Baseline’s 54).
-This is expected from precision tuning: Reval can sample more points by avoiding timeouts, but its overapproximations can also cause it to unnecessarily pass the precision limit.
-In any case, the differences are minimal, showing that precision tuning’s throughput gains come more or less “for free”.
-Sollya, meanwhile, evaluates only 1635 zero values and 52 non-zero values.
-We believe the gap is due to Reval and Baseline’s Rival interval arithmetic library handling overflow and underflow in a more rigorous manner (Flatt and Panchekha, 2023) than Sollya’s MPFI.
-Reval and Baseline also sample 100,346 more points where the correctly-rounded result is infinite; we believe this to be an intentional choice by the Sollya developers, so out of fairness they are not shown in Figure 3.
+Figure 3 compares per-operation execution time by precision for both Reval and Baseline.
+Reval spends significantly less time executing operations at precisions between $2^{6}$ and $2^{11}$, because its non-uniform algorithm assigns these high precisions to only those operations that require them.
+Precision assignment is also fast, as shown by a separate “tuning” column: precision tuning consumes 15.5% of total run time but saves 46.29% of it.
+Figure 4.
+Per-operation precision in final Reval iteration, normalized to maximum precision; left is better.
+A uniform precision would have all weight at 1.0.
+Reval instead evaluates most operations at a dramatically lower than maximum precision, with 53.83% of instructions executed at under 20% of the highest precision.
+This results in its significant performance advantage.
+Figure 5.
+The number of iterations needed for each input in Reval and Baseline; left is better.
+Reval requires fewer iterations than Baseline because it can derive the correct precision instead of searching for it.
+This adds an additional minor performance advantage for Reval.
+The “slack” mechanism is only needed for the second and third iterations, showing it is a minor part of Reval’s algorithm.
+Figure 6.
+The percentage of operations actually evaluated per iteration in Reval; lower is better.
+Reval’s “repeat optimization” allows it to skip re-evaluating an operation if its precision assignment does not change; Baseline’s uniform precision does not allow this optimization so all its data would be at 100%.
+23.77% of operations are skipped on average, adding to Reval’s advantage.
 RQ3
-Figure 4 shows our results for RQ3.
-This plot shows the distribution of execution time by precision across all operations, for both Reval and Baseline.
-(We were not able to log such granular runtime information in Sollya.) Reval saves a significantly fraction of runtime by executing operations at lower precision than Baseline—note the higher bar at precision $2^{6}$ in exchange for lower bars at precisions $2^{7}$ through $2^{11}$.
-Precision assignment is also fast, as shown by the separate column on the right: precision tuning consumes 4.1% of run time but saves 17% of run time.
-Reval spends a tiny fraction of runtime at high precision; in this sense, its precision assignments can be said to be close to optimal.
-Figure 4.
-A weighted histogram showing the distribution of operation execution time by precision, for both Reval and Baseline.
-Lower is better.
-Baseline wastes a substantial fraction of its execution time re-computing the same operations at successively higher precisions, leading to a lot of wasted time at precisions between $2^{7}$ and $2^{11}$.
-By contrast, Reval spends much less time at these high precisions because it can assign high precision only to the few operations that need it.
+Figure 6 shows the main reason behind Reval’s overall better performance.
+This figure shows the non-uniform nature of Reval’s precision assignments, with many operations executed at precisions much lower than the maximum precision assigned: 53.83% at less than 20% of the highest precision.
+By contrast, a uniform precision assignment would have all operations executed at 1.0×1.0$\times$1.0 × the maximum precision.
+Reval’s ability to assign non-uniform precisions allows it to execute most operations at much lower precisions, which leads to its significant performance advantage due to lower data movement, less computation time, and simpler algorithms.
+Figure 6 shows an additional reason behind Reval’s better performance: fewer iterations.
+A standard uniform precision-doubling algorithm must search for the correct precision by trying higher-and-higher precisions.
+Reval instead derives the correct precision mathematically, and thus needs fewer search iterations.
+More precisely, Reval needs only a single tuning iteration for the vast majority (0.73) of inputs, while Baseline needs approximately 3 on average.
+These extra iterations are wasted computation.
+Reval’s “slack” mechanism still performs some search, but this still requires far fewer iterations than Baseline because Reval is still deriving the final precision and also because the slack mechanism’s rarer use means it can be tuned more aggressively.
+For example, after two iterations, 0.24 of Reval’s inputs have converged while only 0.14 of Baseline’s have.
+Figure 6 shows a final, minor reason for Reval’s advantage over Baseline: the repeat optimization.
+Reval’s non-uniform precision assignments often assign the same precisions, iteration to iteration, to an entire subtree, in which case the subtree does not need to be recomputed.
+Baseline’s uniform precision, meanwhile, requires recomputing every operation at every iteration.
+On average, Reval skips 23.77% of instructions, with later iterations skipping slightly more.
+In summary, non-uniform precisions allow Reval to use much lower precisions than Baseline, evaluates them for fewer iterations, and skips many operation evaluations.
+None of these advantages are possible with uniform precisions and add up to Reval’s substantial and growing advantage over Baseline.
+While we weren’t able to gather similarly-granular data about Sollya or another existing real evaluation tool, we expect the same advantages to be present, since all extant tools we know of use the same uniform precision approach.
 Other results
-Besides the core research questions discussed here, Reval’s precision assignment algorithm provides several additional advantages.
-One is robustness: on the approximately 72,450 points that none of the three tools could evaluate, Reval terminates much faster.
-This is because Reval can exit after assigning very high precisions, which takes only one or two iterations.
-For these points, Reval terminates 11.09×11.09$\times$11.09 × faster than Sollya and 19.87×19.87$\times$19.87 × faster than Baseline.
-While these points are not considered in RQ1–RQ3, they play a large role in practical deployments of real evaluation, to the extent that the underlying Rival library uses special flags to detect some kinds of unevaluable inputs (Flatt and Panchekha, 2023).
-Figure 5 breaks down Reval’s greater advantage once unevaluable points are included: early exit allows Reval to skip many useless and expensive high-precision operations, while the uniform precision Baseline has to execute these operations to reach the maximum precision.
-This also makes it practical for Reval to use a significantly higher maximum precision in practical deployments, which in turn allows it to evaluate more inputs.
-Finally, compared to Sollya, Reval supports a broader range of operations (cbrt and atan2 are two examples which show up in the Herbie 2.0 benchmark suite), produces correctly- not faithfully-rounded results, and never returns incorrect results, Sollya produces faithfully-, but not correctly-rounded results in 5,000 cases.
+One additional advantage of Reval is its handling of unevaluable inputs, where it (and Baseline and Sollya) could not evaluate the input due to the maximum precision.
+For these inputs, Reval terminates 6.58×6.58$\times$6.58 × faster than Baseline (and 13.32×13.32$\times$13.32 × faster than Sollya).
+Reval is much faster than Baseline on these inputs because it can typically assign a very high precision after just one or two iterations and then perform an early exit, while Baseline must try every uniform precision before giving up.
+Figure 7 replicates Figure 3 but with these unevaluable inputs included: Baseline now spends a huge amount of time on fruitless evaluations at high precision, while Reval spends very little additional time.
+While these inputs are not considered in RQ1–RQ3, they play a large role in practical deployments, as evidenced by prior work on movability flags to detect some such inputs (Flatt and Panchekha, 2023).
+Because unevaluable inputs can now be swiftly dispatched, non-uniform precision assignments also allow setting a much higher maximum precision; with a uniform precision, raising the maximum precision dramatically raises the cost of unevaluable inputs.
 Figure 7.
 A variant of Figure 3, but over all inputs, including those that neither Reval nor Baseline could evaluate.
-Reval’s early exit allows it to skip most high-precision evaluations, while Baseline wastes a substantial fraction of its runtime uselessly re-evaluating expressions at ever-higher precisions.
+Reval’s early exit optimizations allow it to skip high-precision evaluations of unevaluable inputs, while Baseline must waste time performing.
+This makes Reval more robust to unevaluable inputs and could enable users to set higher maximum precisions.
+Early exit does have some costs: for 47 inputs (of the 567,238 total), Baseline produces a correctly-rounded result while Reval does not.
+This is typically because Reval assigns a precision just over the maximum precision and exits early, while Baseline tries the maximum precision and succeeds.
+This extremely rare occurence can be prevented by enabling Reval’s “lower bound” mode; this allows Reval to evaluate all points that Baseline does, but adds a 0.98% slowdown to Reval.
+Note that this slowdown figure is somewhat deceptive, since it excludes unevaluable points, on which it forces a re-evaluation, at maximum precision, of 18,057 points, causing a significant slow-down.
+As a result, we make this “lower bound” mode off by default: it addresses a very rare issue and hurts robustness to unevaluable inputs.
+Ultimately, we expect Reval users to choose much higher maximum precisions than they would with a uniform precision due to the robustness advantages noted above.
 7.3. Case Studies
-We now examine individual expressions and input points in more detail to highlight aspects of Reval’s precision tuning algorithm.
+We now illustrate three more-complex features of Reval—the slack mechanism, correct rounding, and early exit—by case studies on individual inputs.
 Slack mechanism
-The 2cos benchmark evaluates $f(x,\varepsilon)=\cos(x)-\cos(x+\varepsilon)$ to 53 mantissa bits.
-When $x$ is big and $\varepsilon$ is small, for example $x=10^{300}$ and $\varepsilon=10^{-300}$, the terms being subtracted differ by a tiny amount, making the output hard to compute.
-To evaluate the output correctly, the $x+\varepsilon$ term must be computed to at least 600 digits, or about 2000 bits.
-Then the $\cos$ operations have to be evaluated to approximately 300 digits, or about 1000 bits, and finally the subtraction needs to be evaluated to only 53 bits to achieve a correctly-rounded result.
-Reval achieves this precision assignment in three iterations.
-In the first, untuned iteration, Reval is able to estimate the sizes of $x$, $x+\varepsilon$, and $\cos(x)$.
-However, because this first iteration uses low precision, $x+e$ covers more than one full period of $\cos$—in fact, it covers approximately $3\cdot 10^{282}$ of them.
-The interval computed for $\cos(x+e)$ is thus $[-1,1]$, and the subtraction operation likewise has a wide computed interval crossing zero.
-As a result, during the first tuning pass, Reval applies the slack mechanism twice, once to the subtraction and once to cosine.
-This results in a precision assignment of $59073+5+\operatorname{slack}[0]=$ 590 for both $\cos$ operations, the slack being introduced by the subtraction, and $590+5+1000+\operatorname{slack}[0]\approx 2100$ for $x+\varepsilon$, where $590$ is the output precision needed, $1000$ is the amplification factor for cosine, and the slack being introduced by the cosine.
-The amplification factor of $1000$ is a good demonstration of how slack differs from naive uniform precisions.
-Even though the slack mechanism is used, the $\cos$ operation still provides some information.
-Specifically, the input to $\cos$ has to be precise enough to at least identify what period of $\cos$ it is in.
-The full precision assignment is shown in Figure 6.
-The first Reval precision assignment:
-$\begin{array}[]{llll}P[-]&&&=73\\ P[\cos(x)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil&=73+5+\operatorname{slack}[0]&=590\\ P[\cos(+)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{2}(-)\rceil&=73+5+\operatorname{slack}[0]&=590\\ P[+]&=P[\cos(+)]+5+\lceil\log_{2}\operatorname{ampl}_{1}(\cos(+))\rceil&=590+5+1000+\operatorname{slack}[0]&=2107\end{array}$
-The second Reval precision assignment:
-$\begin{array}[]{llll}P[-]&&&=73\\ P[\cos(x)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil&=73+5+593+\operatorname{slack}[1]&=1695\\ P[\cos(+)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{2}(-)\rceil&=73+5+593+\operatorname{slack}[1]&=1695\\ P[+]&=P[\cos(+)]+5+\lceil\log_{2}\operatorname{ampl}_{1}(\cos(+))\rceil&=1695+5+1000&=2700\\ \end{array}$
+$\begin{array}[]{llllll}z_{i}&\text{First tuning iteration}&\text{Second tuningiteration}\\ \hline\cr z_{4}=z_{3}-z_{2}&53+10=63&53+10=63\\ z_{3}=\cos(x)&63+5+\operatorname{slack}[0]=580&63+5+(593+\operatorname{slack}[1])=1685\\ z_{2}=\cos(z_{1})&63+5+\operatorname{slack}[0]=580&63+5+(593+\operatorname{slack}[1])=1685\\ z_{1}=x+\varepsilon&580+5+(1000+\operatorname{slack}[0])=2097&1685+5+1000=2690\\ \end{array}$
 Figure 8.
 Precision assignment in the first and second iteration for the slack mechanism case study.
-Note that, even though the slack mechanism partially re-introduces precision doubling, amplification factors still provide useful information.
+Note that, even though the slack mechanism partially re-introduces search, amplification factors still provide useful information and Reval still converges to a high precision much more quickly than Baseline.
+Consider the 2cos benchmark, $\cos(x)-\cos(x+\varepsilon)$, with inputs $x=10^{300}$ and $\varepsilon=10^{-300}$.
+[^22] This input (and others in the case studies) are a simplified form of one of the 567,238 inputs.
+The terms being subtracted differ by a tiny amount, and need to be evaluated to approximately 300 digits (1000 bits) to achieve a correctly-rounded result.
+This in turn requires evaluating $x+\varepsilon$ to at least 600 digits, or about 2000 bits.
+Reval achieves this precision assignment in two tuning iterations, shown in Figure 8.
+After the first, untuned evaluation, Reval is able to estimate the sizes of $x$, $x+\varepsilon$, and $\cos(x)$.
+However, because this first iteration uses low precision, $x+e$ covers a wide interval and the interval computed for $\cos(x+e)$ is thus $[-1,1]$, and the subtraction likewise has a wide interval crossing zero.
+As a result, during the first tuning pass, Reval applies the slack mechanism twice, once to the subtraction and once to cosine, ultimately assigning a precision of $58063+5+\operatorname{slack}[0]=$ 580 for both $\cos$ operations, the slack being introduced by the subtraction, and $2097580+5+1000+\operatorname{slack}[0]=$ 2097 for $x+\varepsilon$, where $1000$ is the amplification factor for cosine, and the slack is introduced by the cosine.
+This amplification factor of $1000$ is a good demonstration of how slack differs from naive uniform precisions: even though the slack mechanism is used, the $\cos$ operation still provides some information and allows Reval to make a more useful guess.
 After re-evaluating with these new precision assignments, the output of the $\cos$ operation no longer crosses zero, but the output of the subtraction still does.
-The second precision assignment, then, uses the slack mechanism only for the subtraction, assigning a precision of $169573+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil=78+593+\operatorname{slack}[1]=$ 1695 to both $\cos$ operations and a precision of $1695+5+1000=2700$ to the $x+\varepsilon$ term.
+The second precision assignment, then, uses the slack mechanism only for the subtraction, assigning a precision of $168563+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil=68+593+\operatorname{slack}[1]=$ 1685 to both $\cos$ operations and a precision of $1685+5+1000=2690$ to the $x+\varepsilon$ term.
 A third re-evaluation then results in a correctly-rounded result.
-Note that, while Reval needed only 3 evaluations, a uniform-precision evaluation would require 5 iterations.
+Note that, while Reval needed three evaluations, a uniform-precision evaluation would require evaluations at $2^{6}$, $2^{7}$, $2^{8}$, $2^{9}$, $2^{10}$, and $2^{11}$ before finally succeeding at a precision of $2^{12}$.
 Correct rounding mechanism
-The placeG benchmark evaluates $f(x,y,z)=(x+y)\times(z+1)$ to 53 mantissa bits.
-Since the benchmark uses only multiplications and divisions, the exact real result can be represented in floating-point.
-Consider positive $x$, $y$, and $z$; then all the $\operatorname{ampl}$ and $\operatorname{intro}$ factors are 1 or less.
-Reval thus assigns low precisions to all operations: 73 bits to the multiplication and 78 to both additions.
-That works for many positive inputs, but not all.
-Some specific inputs suffer from double rounding, such as $x=1.3002052657264033\times 10^{189}$, $y=3.084776002356433\times 10^{188}$, and $z=1\times 10^{-200}$ in double-precision floating-point.
-These $x$ and $y$ values are exactly one exponent apart, and both end in a 1 bit.
-This means that their sum, $x+y$, is exactly 54 bits wide, meaning it lies exactly on a rounding boundary.
-That means it rounds up if $1+z$ is greater than $1$, and down (toward even) if $1+z$ is equal to $1$.
-In fact, $z+1$ is very slightly greater than 1, which means the answer should round up, but because $z$ is much smaller than $1$, $1+z$ needs to be computed to many, many bits to conclude that it is in fact greater than $1$.
-Specifically, in 78 bits, $z+1$ is computed to be $[1,1.000000000000000000000007]$ in the first iteration and the overall expression is computed to be $[1.6...\times 10^{189},1.6...\times 10^{189}]$, where the two endpoints of that interval are successive floating-point numbers.
-This triggers the double rounding mechanism, causing Reval to evaluate the expression first to $58573+\operatorname{slack}[0]=$ 585 and then to 73+slack[$109773+\operatorname{slack}[1]=$ 1097 bits.
-This last iteration finally evaluates $1+z$ exactly, proving that it is strictly greater than $1$.
-By increasing the overall precision demands, the double rounding mechanism was able to eventually determine which side of the rounding boundary the result was on.
+Consider the placeG benchmark, $(x+y)\times(z+1)$, at the double-precision inputs $x=1.3002052657264033\times 10^{189}$, $y=3.084776002356433\times 10^{188}$, and $z=1\times 10^{-200}$.
+Since all inputs are positive and the only operations are addition and multiplication, all $\operatorname{ampl}$ and $\operatorname{intro}$ factors are 1 or less and Reval thus assigns 64 bits of precision to all operations.
+However, these precisely-chosen $x$ and $y$ values are exactly one exponent apart and both end in a 1 bit.
+This means that their sum, $x+y$, lies exactly on a rounding boundary, rounding up if $z+1>1$ and down otherwise.
+This triggers the double rounding mechanism, causing Reval to evaluate the expression first to $57563+\operatorname{slack}[0]=$ 575 and then to 63+slack[$108763+\operatorname{slack}[1]=$ 1087 bits.
+Only after this second re-evaluation can Reval prove that $z+1$ is strictly greater than $1$ and achieve a correctly-rounded result.
 Early stopping
-The 2nthrt benchmark evaluates $f(x,n)=(x+1)^{1/n}-x^{1/n}$ to double precision.
-The benchmark is hard to evaluate for large $x$ and small $n$—say, $x=10^{200}$ and $n=10^{-200}$.
-These inputs require raising a large number to a large power, causing overflow: the intermediate result of $(x+1)^{1/n}$ is $[\approx 2\times 10^{323,228,496},+\infty]$ after the first iteration.
+Consider the 2nthrt benchmark, $(x+1)^{1/n}-x^{1/n}$, for $x=10^{200}$ and $n=10^{-200}$.
+These inputs require raising a large number to a large power, causing overflow: $(x+1)^{1/n}$ evaluates to $[\approx 2\times 10^{323,228,496},+\infty]$ after the first, untuned evaluation.
 The $x^{1/n}$ term likewise overflows, resulting in an output interval of [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] after subtracting the two.
-When adjusting precisions for the second iteration, precision tuning algorithm calculates an amplification factor for the subtraction of
+When adjusting precisions for the second iteration, the amplification factor for the subtraction is
 […]
-where each term contributes a slack term.
-This causes Reval to assign a precision of billions of bits to the left-hand term, higher than the maximum allowed precision 32 256.
-Reval therefore terminates evaluation early.
+This causes Reval to assign a precision of billions of bits to the left-hand term, higher than the maximum allowed precision 32 256. Reval therefore, correctly, terminates evaluation early.
 8. Related Work
@@
 Early implementations used lazy digit sequences, but later work (Lee and Boehm, 1990; Boehm et al., 1986) showed that Ziv’s strategy (Ziv, 1991) with interval arithmetic is typically faster.
-Interval arithmetic is an old and widely-used technique.
-Prominent implementations include Arb (Johansson, 2016), Moore (Mascarenhas, 2016), MPFI (Revol and Rouillier, 2005), Rival (Flatt and Panchekha, 2023), and others.
-Interval arithmetic is used in the Sollya environment (Chevillard et al., 2010) for numerical programming and possibly in Wolfram Mathematica (Flatt and Panchekha, 2023).
+Prominent interval arithmetic implementations include Arb (Johansson, 2016), Moore (Mascarenhas, 2016), MPFI (Revol and Rouillier, 2005), Rival (Flatt and Panchekha, 2023), and others.
 Many researchers have proposed various tweaks to computable real implementations, such as movability flags to track overflow (Flatt and Panchekha, 2023).
 Perhaps the most similar to the current work was proposed by Li and Yong (2007): mixed-precision evaluation using “weights”.
-However, they seem to expect the weights to be user-controlled, whereas Reval derives the precision tuning automatically.
+However, this work expects users to manually set weights, whereas Reval derives a non-uniform precision assignment automatically.
 Interval arithmetic and constructive reals are closely related to research into automatic worst-case error bounds for floating-point programs.
@@
 While this provides an algorithmic speedup, ATOMU did not evaluate it; instead, ATOMU focused on condition numbers as a method for driving input generation.
-In most worst-case error bound tools, the goal is achieving the lowest possible sound error bound, unlike in this paper, where the focus is primarily on computational efficiency.
+In most worst-case error bound tools, the goal is achieving the smallest possible sound error bound, unlike in this paper, where the focus is primarily on computational efficiency.
 Automatic precision tuning is also an active area of research, starting with Precimonious (Rubio-González et al., 2013), which used delta-debugging to assign precisions while evaluating accuracy on representative inputs.
@@
 9. Conclusion
-Evaluating real expressions to high precision requires using high-precision interval arithmetic.
+Evaluating real expressions to high precision is a critical component of algorithms across science, computing, and especially numerics.
 But state-of-the-art tools use a uniform precision, doubled with every iteration.
@@
 A key optimization is the use of “exponent tricks”, which allow computing amplification factors using only operations on machine integers.
-As a result, Reval is able to evaluate real expressions 1.25×1.25$\times$1.25 × faster than the state-of-the-art Sollya tool across our suite of 528 benchmarks, with Reval’s advantage rising to 2.99×2.$9$ for the most difficult input points.
+As a result, Reval is able to evaluate real expressions 1.47×1.47$\times$1.47 × faster than the state-of-the-art Sollya tool across our suite of 493 benchmarks, with Reval’s advantage rising to 4.92×4.92$\times$4.92 × for the most difficult input points.
 It achieves this speedup largely by using a sophisticated mixed-precision assignment algorithm that evaluates operations at high precision only when necessary.